### v1

# Modeling Innovation Ecosystem Dynamics through Interacting Reinforced Bernoulli Processes

Abstract.
Understanding how capabilities evolve into core capabilities—and how core capabilities may ossify into rigidities—is central to innovation strategy (Leonard-Barton 1992, Teece 2009). A major challenge in formalizing this process lies in the interactive nature of innovation: successes in one domain often reshape others, endogenizing specialization and complicating isolated modeling. This is especially true in ecosystems where firm capabilities and innovation outcomes hinge on managing interdependencies and complementarities (Jacobides et al. 2018, 2024).
To address this, we propose a novel formal model based on interacting reinforced Bernoulli processes. This framework captures how patent successes propagate across technological categories and how these categories co-evolve. The model is able to jointly account for several stylized facts in the empirical innovation literature, including sublinear success growth (success-probability decay), convergence of success shares across fields, and diminishing cross-category correlations over time.
Empirical validation using GLOBAL PATSTAT (1980–2018) supports the theoretical predictions. We estimate the structural parameters of the interaction matrix and we also propose a statistical procedure to make inference on the intensity of cross-category interactions under the mean-field assumption.
By endogenizing technological specialization, our model provides a strategic tool for policymakers and managers, supporting decision-making in complex, co-evolving innovation ecosystems—where targeted interventions can produce systemic effects, influencing competitive trajectories and shaping long-term patterns of specialization.
Key-words: Innovation Ecosystems, Dynamic Capabilities, Interacting Bernoulli Processes, Patent Interactions, Technological Specialization
1. Introduction
Understanding the extent to which capabilities evolve into core capabilities—and how core capabilities may turn into core rigidities (Leonard-Barton 1992)—is a crucial question for firms, particularly those that are innovation-driven (Teece 2009, Eisenhardt and Martin 2000, Helfat and Peteraf 2003). Salter and Alexy (2014) support this view by emphasizing that a well-balanced mix of skills within teams is a fundamental driver of innovation—particularly when these skills are leveraged within a proactive ecosystem capable of transforming them into core capabilities (Jacobides et al. 2018). Finding such a mix of skills within an ecosystem is, however not an easy task and further exploration is needed. An interesting hint in this direction is provided in Kim and Magee (2017) and in Pichler et al. (2020). They demonstrate that knowledge flows predominantly occur within previously linked technological domains, which implies that domain-specific learning can reinforce the development of specialized skill sets (Cohen et al. 1990, Malerba and Orsenigo 1997). In their study, as well as in several others (Teece et al. 1997, Adner and Kapoor 2010, Youn et al. 2015, Pugliese et al. 2019) it clearly emerges how the concept of skills and capabilities is inherently linked with the concept of technological fields. Studying how technologies, inherently linked to specific skill requirements, interact within an ecosystem is, therefore, crucial for formalizing the transformation of these skills into core capabilities (Granstrand 1998, Kim et al. 2022). One of the main challenges posed by the interaction among innovations is that they endogenize the processes of specialization and the formation of core capabilities. In other words, the dynamic and path-dependent nature of capability evolution (Teece 2009, Eisenhardt and Martin 2000) is shaped by the co-evolution of technologies and markets. As a result, the interplay among innovations does more than merely shifting the external environment—it recursively shapes the internal structure and evolutionary trajectory of firm capabilities, making it difficult to model them in isolation. This challenge is well recognized in the literature, particularly given that firm capabilities operate within ecosystems whose value creation depends on how multilateral interdependencies–based on various forms of complementarities–are managed (Jacobides et al. 2018). Both generic and idiosyncratic complementarities across capabilities– and thus across technological domains– are crucial determinants of a firm’s success in innovation (Jacobides et al. 2024). An example of this is given by innovation-based platforms, where developers must be incentivized to generate complementary innovations (Jacobides et al. 2024).
Given the above premises, formalizing the processes that govern interactions among innovations across different technological domains becomes essential for determining the extent to which ecosystems can effectively exploit complementarities—thereby fostering the development of core capabilities while avoiding their degeneration into core rigidities. The aim of the present work is to develop a novel mathematical framework to rigorously formalize such processes of interactive innovations.
Technological innovation is not merely a process of isolated breakthroughs but is strongly influenced by self-reinforcing structures, where mutually connected technological fields benefit from each other, leading to cumulative innovation dynamics (Salter and Alexy 2014, Youn et al. 2015, Napolitano et al. 2018, Castaldi et al. 2015). The relationship between cross-category interactions and innovation success is, however, not necessarily straightforward and necessitates a dedicated formal treatment. While interactions within closely related Cooperative Patent Classification (CPC) categories may increase the overall likelihood of incremental innovation, interactions across technological domains may be a key driver of breakthrough innovations—though such events remain rare (Castaldi et al. 2015). Kovács et al. (2021), for instance, show that patents classified in high-contrast categories—those with clear boundaries from neighboring technologies—receive more attention and citations than those in low-contrast categories. This suggests that classification is not merely an administrative process but actively shapes innovation diffusion by influencing how patents are discovered, cited, and integrated into future technological developments. According to Kovács et al. (2021), one key mechanism through which this occurs is search dynamics—the way patent users identify and engage with prior knowledge. Their findings reinforce the idea that the way technologies are classified and interconnected affects not only their immediate visibility but also their long-term influence on the technological evolution of firms, seen as innovation ecosystems (Coccia 2017, Colladon et al. 2025). Specifically, understanding these systemic interdependencies in technologies is central to studying the logic of search dynamics and, consequently, to modeling how capabilities evolve within firms through a clearer understanding of their technological evolution (Pichler et al. 2020).
The literature on this topic has primarily documented the clustering of technological domains and their reinforcing interactions (Youn et al. 2015), yet it often falls short in explicating the underlying mechanisms—highlighting the need for a formal model. Indeed, while Leonard-Barton (1992) emphasized that core capabilities may evolve into core rigidities if not continuously adapted, numerous scholars over the years have argued that a formal framework is indispensable for meaningfully quantifying and understanding this transformation (Teece 2009, Zollo and Winter 2002, Helfat and Peteraf 2003, Youn et al. 2015). The latter should model the recursive, path-dependent interactions among innovations and pave the way for future research to assess how these interactions influence the development, entrenchment, or renewal of capabilities. Our study bridges this gap by introducing a mathematical framework to assess how innovation success in one CPC category reinforces success, not only in the same category, but also in the other categories, thus offering a structured explanation for how technological trajectories emerge and evolve. A major difficulty in building such a formal framework lies in the fact that, despite the growing recognition of cross-category knowledge flows (Blazsek and Escribano 2010, Nemet 2012, Pichler et al. 2020, Singh et al. 2021), the role of cross-category interactions in innovation remains largely unexplored in formal models.
The relevance of our model and its ability to explain real-world innovation dynamics is assessed by identifying a set of stylized facts some of which have been consistently documented in the empirical literature. While existing models typically focus on a subset of these phenomena, our framework provides a unified theoretical explanation that links them through the structure of interactions among technological domains as required by the literature (Pichler et al. 2020). In few words, our model consists of a finite set of interacting Bernoulli processes, each for each technological category. The time-steps are marked by the succession of patents, according to their full date of publications. The status $1$ of the process related to category $h$ identifies a patent which is a “success”, with respect to a novelty measure, in the technological field $h$. The dynamics of the Bernoulli processes are governed by a self- and cross-category reinforcement rule: the probability that patent $n+1$ will be a success for category $h$ has an increasing dependence on the number of past successes, not only in category $h$, but also in all the other categories. The interaction matrix $\Gamma=(\gamma_{j,h})_{j,h}$ collects the weight of this dependence: $\gamma_{j,h}\geq 0$ quantifies how much past successes in a given category $j$ induces future successes in category $h$. As a novelty measure for our model, we use a variant of the forward citation index in Squicciarini et al. (2013): essentially, in order to measure the impact of a given patent in a certain CPC-1 category, we measure the number of citations to the considered patent by subsequent patents belonging to the considered category, normalized with respect to the patent cohort. However, it is very important to note that our methodology is presented in general so that it can be applied with different classifications in categories and/or different novelty measures, even in other contexts. [^1] The interested reader is referred to Section A in appendix for a discussion on novelty measures. There we also motivate our preference for the index of Squicciarini et al. (2013) over other measures of patent novelty.
All the provided theoretical results are proven rigorously and are supported by the analysis of a real-world data set from the GLOBAL PATSTAT database. More specifically, we analyze all the patents published in the period $[1980-2018]$ showing that the observed behaviors match with the theoretical ones. We also perform a statistical study on the interaction intensity among the categories.
Below, we present a list of core empirical regularities that our model is able to jointly explain, offering a more comprehensive perspective on the nature of technological path dependence. We divided the mentioned regularities into those related to the formation of core capabilities—denoted as $\mathcal{C}$.i—and those related to the ossification of core capabilities into core rigidities—denoted as $\mathcal{R}$.i.
  • $\mathcal{C}$1
    Early-stage interventions produce stronger and more persistent effects, especially before the innovation landscape fragments into relatively autonomous and specialized domains. [^2] Similar calls for timely and anticipatory policy actions have been raised in studies emphasizing the cumulative and path-dependent nature of technological change (David 1985, Mazzucato 2018). Our model is ruled by a self- and cross- reinforcement mechanism so that the probability of a future success in a certain category depends on the number of past successes in that category and also in all the other categories. This study therefore highlights the importance of early interventions as having the strongest systemic impact, thereby reinforcing Leonard-Barton (1992)’s view that the evolution of core capabilities is path-dependent and resistant to late-stage adjustments—emphasizing the strategic imperative for timely, anticipatory actions in innovation policy and capability building.
  • $\mathcal{C}$2
    The probability of observing a future success decreases along time. This fact is reflected in a sub-linear growth of the number of observed successes: indeed, for each category $h$, the number $S_{t,h}$ of successes in category $h$ exhibits a power-law growth with an Heaps’ exponent lower than one. In other words, technological categories become ”crowded” over time, with increasing saturation and depletion of easy innovation opportunities, making new breakthroughs less frequent (Jones 2009, Bloom et al. 2020, Clancy 2023). In a broader innovation ecosystem, this supports the need for cross-category recombination: when a single field is exhausted, future breakthroughs often require combining insights from other fields (Fleming 2001), thereby creating new opportunities of developing core capabilities.
  • $\mathcal{C}$3
    Irreducibility of the interaction matrix: all the categories present the same Heaps exponent and this means that all the categories interact each other or directly (when $\gamma_{j,h}>0$) or undirectly by means of a path in the graph associated to the interaction matrix $\Gamma$. This evidence formalizes the arguments in Weitzman (1998), Fleming (2001), Alves et al. (2007), Godin (2017) among others.
  • $\mathcal{R}$1
    Relative number of successes across categories stabilizes along time: for two different categories $h$ and $j$, the ratio $S_{t,h}/S_{t,j}$ stabilizes over time toward the ratio of their respective components of the eigenvector-centrality score vector ${\mathbf{u}}$, that is $S_{t,h}/S_{t,j}\to u_{h}/u_{j}$. This means that the relative proportions of successes across categories remain stable in the long run. Such stability reflects the consolidation of core capabilities within firms constituting the innovation ecosystem, on the one hand paving the way for the formation of new core capabilities and on the other starting the process of ossification of traditional ones into core rigidities (Leonard-Barton 1992). While related ideas regarding the influence of central nodes over time in economic production networks have been theorized (Acemoglu et al. 2016), to the best of our knowledge, this specific idea has never been formally stated or formalized within the innovation literature. [^3] Although empirical evidence point in this direction (Breschi et al. 2000, Malerba 2002).
  • $\mathcal{R}$2
    The long-run decline of the cross-category correlations: if we denote by $X_{n,h}$ the random variable that takes the value $1$ if patent $n$ is a success in category $h$ and 00 otherwise, then the correlation coefficients between $X_{n,h}$ and $X_{n,j}$ decreases to zero along time. Indeed, although the interaction matrix $\Gamma$ is constant along time, its impact on the probabilities of success decreases linearly along time, generating vanishing correlations among the categories. This last piece of evidence is reminiscent of the knowledge spillovers literature (Jaffe and Trajtenberg 2002). In their book (and subsequent works), Jaffe and Trajtenberg (2002) observe increasing specialization within technological fields, implying that cross-field citations — and, by extension, indirect correlations — tend to diminish over time relative to within-field citations. However, the treatment in Jaffe and Trajtenberg (2002) remains primarily empirical and descriptive. This process of specialization is at the basis of the ossification process into core rigidities described in Leonard-Barton (1992).
A more thorough positioning of our contribution within the existing literature, considering the empirical regularities discussed above, is presented in Section 1.1.
Our novel mathematical model offers a strategic tool for policymakers and practitioners. Indeed, as aforementioned, it helps managers understanding how and when innovation processes interact, when a technological field becomes saturated, and how it relates to other fields over the long term. This, in turn, provides valuable insights to identify how and when the capabilities—required for developing innovations within a given technological field—may evolve into core capabilities or degenerate into core rigidities (Leonard-Barton 1992).
Furthermore, our formal model can guide managers in innovation-related decision-making, which has become more challenging due to the growing complexity and variety of innovation ecosystems (Dias Sant´ Ana et al. 2020), as well as the increasing interdependencies among technologies (Adner and Kapoor 2010, Russell and Smorodinskaya 2018, Granstrand and Holgersson 2020) –especially in an environment where innovation ecosystems coevolve (Thomas and Autio 2019, Paasi et al. 2023). In particular, we illustrate how a localized innovation shock in one technological domain can propagate across the broader ecosystem, influencing success rates in other domains depending on the structure of interdependencies. Importantly, we show that even targeted interventions can have far-reaching effects—either amplifying or constraining innovation elsewhere—thereby offering managers additional tools to assess competitive dynamics and make strategic decisions within innovation ecosystems.
In the next subsection we position our work in relation to the existing literature.
1.1. Positioning within the literature
Prior studies in the management literature have generally addressed aspects of patent dynamics—such as citation behavior (Lee et al. 2012), innovation competition (Keller 2007), and R&D investment (Akcigit and Kerr 2018, Aghion and Howitt 2007)— lacking, however, a formalization of the process governing cross-category technological interactions. In the course of the years scholars have deepen the latter aspect by investigating knowledge spillovers across technologies (Kim and Magee 2017) and forming patents’ networks (Higham et al. 2022). Emerging methods, including machine learning models for predicting patent success (Yang 2023), and descriptive studies on radical inventions’ cross-domain influence (Dahlin and Behrens 2005), highlight this gap. These works, however, remain empirical in nature and do not provide a formal understanding of the processes that drive their results. The present paper addresses this by explicitly capturing – via interactive Bernoulli processes– how innovation in one CPC category affects outcomes in others, providing a structural model of technological co-evolution.
Parallel developments in management research, often leveraging NLP and LLMs, have focused on patent similarity across categories (Younge and Kuhn 2016, Arts et al. 2018, Jurek 2024, Bekamiri et al. 2024). While useful for mapping static proximity, these approaches remain correlational and fall short in explaining dynamic mechanisms such as technological convergence, divergence, and the long-run distribution of innovation outcomes (Ganguli et al. 2024). Even studies incorporating dynamics (Kim and Magee 2017, Park and Magee 2017, Taalbi 2023, Maragakis et al. 2023) tend to remain empirical, lacking formal structure.
As innovation increasingly unfolds through technological interdependencies over time – generative AI constitutes an example of this phenomenon (Chen et al. 2023, Du et al. 2024, Mohammadabadi 2025)– recent works emphasize the need for formal models that endogenize these relationships (Pichler et al. 2020, Taalbi 2023). Existing models, in particular, are often capable of showing isolated phenomena but do not jointly address long-run specialization, capability evolution, or structural path dependence. Our contribution lies in bridging this gap by introducing a unified framework that models interacting technological domains, capturing features such as long-run dominance, category cycles, and the evolution from interdisciplinarity to specialization (Wang and Schneider 2020).
We now consider the main features of our model with respect to the literature:
  • Early-stage interventions produce stronger and more persistent effects. Back at the time David (1985) introduced the concept of path dependence in technological change, illustrating how early choices can lock-in trajectories that resist later correction. This inertia can either evolve into rigidities if not shaped proactively (Leonard-Barton 1992), or be catalyzed into core capabilities through well-designed, mission-oriented innovation policies implemented by institutions (Mazzucato 2018, Kattel and Mazzucato 2018). Our paper elucidates these dynamics by examining how innovation processes interact with one another, thereby potentially influencing the early-stage outcomes of adjacent processes.
  • The probability of observing a future success decreases along time. This aspect is reflected in a power-law growth of the number of observed successes with an exponent smaller than one. This evidence was noted in Heaps (1978) (indeed, the power-law exponent is often called the Heaps’ exponent) and observed in many fields, such as economics, natural sciences, linguistics and so on. Since Heaps’ work, it has become natural to require that applicative models satisfy this power-law property (Clauset et al. 2009, Mitzenmacher 2004, and 2005).
  • All categories interact with one another, either directly or indirectly. Several studies in the literature have hypothesized or empirically observed this pattern (Verspagen 2007, Rigby 2015, Higham et al. 2022). Few, however have provided a formal modelization of this phenomenon. We did so by introducing an interaction matrix governing the probability of future success in a certain category.
  • Relative number of successes across categories stabilizes along time. Despite continuous innovation, patent categories exhibit long-term stability in their relative success rates. We demonstrate that the ratio of cumulative successes across categories converges to the ratio of their centrality scores, indicating a predictable long-run distribution. This fact aligns with the empirical observation that technological domains rarely disappear entirely (Park and Magee 2017, Kim and Magee 2017) and an interpretation can be given in light of the work of Maslach (2016): firms persist after incremental failures but pivot after radical ones. As most innovation is incremental in nature, firms tend to remain anchored in familiar technological domains, reinforcing their sustained importance despite the emergence of new fields. [^4] This also aligns with empirical observations that technological domains, such as pharmaceuticals and automotive, maintain consistent importance over time (Maslach 2016, Perri et al. 2020, Lafond and Kim 2019)
  • The long-run decline of the cross-category correlations. Our model captures the declining correlation between patent categories over time, reflecting a trend toward increased specialization –a well-documented phenomenon in the literature (Ganguli et al. 2024). Namely, technologies that were initially interdisciplinary tend to become more specialized as they mature. A well-documented historical example of this phenomenon is the divergence between telecommunications and information technology (IT) (Messerschmitt 1996, Shin 2010). [^5] Initially, these domains were tightly interconnected, with innovations in data transmission and computing co-evolving (Jones and McGuirt 1991, Fransman 2001, Leiner et al. 2009). Over time, however, they have become increasingly specialized: telecommunications has focused on network infrastructure and communication protocols, while IT has expanded into computing systems, software, and data management (Messerschmitt 1996, Garrone et al. 2002).
Finally, a clarification is in order: the literature has adopted diverse definitions of the terms novelty, innovation, and invention, often depending on the degree of novelty introduced or the social and economic impact of the innovation. In physics and mathematics, innovation processes are typically modeled as sequences of observed elements. An item is labeled a novelty or innovation if it appears for the first time in a sequence of observations; otherwise, it is classified as a repetition of a specific element observed in the past. A classic example is provided by the analysis of word sequences in texts, as explored by Tria et al. (2014), where a novelty is defined as the first occurrence of a word in the sequence, and subsequent appearances of the same word are treated as repetitions. In this framework, suitable models are the Poisson-Dirichlet model (Pitman and Yor 1997) and the urn with triggering (Tria et al. 2014) for a single process and their corresponding versions with interaction (Fortini et al. 2018, Iacopini et al. 2020, Aletti et al. 2023a, 2025).
In contrast, in economics, innovation is frequently synonymous of patented invention. Here, a patent is, by definition, a novel item—never previously observed—but the emphasis is on whether the patent constitutes a successful innovation. Economic studies thus often model the probability of generating a significant innovation—what we refer to in this work as a “success”. Accordingly, in this paper, we use the terms innovation, novelty, and invention interchangeably, focusing primarily on whether a patented item meets the threshold of success (with respect to some measure), regardless of whether it constitutes a breakthrough or an incremental improvement. Moreover, in this framework, the natural class of models are the interacting Bernoulli processes, which our model belong to.
To the best of our knowledge, there are only few papers in the scientific literature regarding interacting (or interrelated) Bernoulli processes. The first one is Antelman (1972), which tries to develop a Bayesian method for making statistical inference on the parameters of two Bernoulli processes when the a-priori joint distribution that generates them exhibits dependence. More precisely, a family of closed-under sampling prior distributions, called the Dirichlet-beta family are introduced. However, these distributions result intractable and several approximation are necessarily proposed. Moreover, this study does not cover the case of more than two Bernoulli processes, as in our case. Another paper is Katselis et al. (2019), where the parameters of the Bernoulli distributions at time $t+1$ depend only on the status of the processes at the present time $t$ so that a multi-lag dependence is not allowed. Finally, in Pandit et al. (2019) the provided model relates the parameter of the Bernoulli distribution of a certain process at a given time-step to the history of all the processes over $L$ time lags by means of interaction parameters $\theta_{h,j,\ell}$, that capture how much process $j$ at time lag $\ell$ affects process $h$. Hence, this model and ours share the dependence of the parameter of the Bernoulli distribution for the future observations on the present and past observations of all the processes. However, important differences are present. First of all, in Pandit et al. (2019) the dependence on the past is up to a fixed number $L$ of lags, while in our model the dependence is on all the past observations. Moreover, in Pandit et al. (2019) the parameters $\theta_{h,j,\ell}$ do not depend on the time-step $t$, while in our model the corresponding quantities essentially are $\gamma_{j,h}/(c+t)$, that decreases along time and let us to get, for each process, a vanishing success probability and a power-law behavior of the number of successes, that we observe in the patent data set. Finally, the paper Pandit et al. (2019) focuses on a maximum likelihood estimation procedure for the model parameters $\theta_{h,j,\ell}\in{\mathbb{R}}^{N\times N\times L}$, that, since the high number of parameters, needs some sparsity condition on $\Theta=\{\theta_{h,j,\ell}\}$; while we provide a deep theoretical study of the model obtaining asymptotic results that can be verified on real data.
It is also worthwhile to stress that the theoretical results for our model follow from new general theoretical results, that, among other things, have the merit to provide a general mathematical framework where we can also include the results in the above quoted papers Tria et al. (2014), Aletti et al. (2023b) regarding the number of distinct observed novelties.
1.2. Structure of the paper
The remainder of the paper is organized as follows. In Section 2, we introduce our primary novelty measure based on a category-level variant of the forward citation index, and define what constitutes a “success” in each technological field. Section 3 presents the proposed mathematical model of interacting Bernoulli processes, offering a theoretical framework to capture the dynamics of innovation across categories, and states the main asymptotic results. All the theoretical results are analytically proven in the appendix (supplementary material). Moreover, Section 3 highlights the implications of the model through a simulation exercise, related to the propagation effects of exogenous shocks. In Section 4, we validate the theoretical findings using a comprehensive dataset of patents from the GLOBAL PATSTAT database, and we estimate the structural parameters of the interaction matrix. Statistical inference on the interaction intensity under a mean-field assumption is also performed. Section 5 concludes.
2. Category forward citation index and definition of “success”
As previously noted in Subsection 1.1, in economics the term “innovation” often refers to “patented invention”. Since patents are novel by definition, our focus lies on whether they represent a successful innovation. Thus, we use the terms innovation, novelty, and invention interchangeably, emphasizing their success rather than their degree of novelty.
The number of citations a given patent receives (more briefly, forward citations) reflects the technological importance of the patent for the development of subsequent technologies (see Squicciarini et al. (2013) and reference therein). In other words, forward citations to a certain patent represent the range of later inventions that have benefited from that patent. Therefore, we have decided to use a variant of the forward citation index in Squicciarini et al. (2013) for detecting the patents that could be defined a “success” in one or more categories. However, it is worthwhile to note that our approach is not specific for a certain classification in categories or for the chosen index and so it could be also applied with a different index aimed to evaluate the importance of a patent. [^6] A review of the novelty indices present in the literature and potentially applicable to our context is presented in Appendix A. There we also motivate our preference for the index of Squicciarini et al. (2013) over other measures of patent novelty.
Consider a system $\mathcal{H}$ of technological categories for the classification of patents. For a given patent $n$, with publication (full) date $d_{n}$, we set, for each category $h\in{\mathcal{H}}$,
$CIT_{n,h}$= number of patents with publication date in the period $[d_{n},d_{n}+365\times T]$ and belonging to category $h$ and citing $n$.
Then, given a patent $n$ with publication year $y_{n}$ and belonging to category $c_{n}$, we refer to the set of patents with the same publication year of $n$, i.e. $y_{n}$, and belonging to the same category of $n$, i.e. $c_{n}$, as the “cohort of patent $n$. Hence, for each category $h$, we define the index $I_{n,h}$ as the normalization, with respect to the cohort of patent $n$, of the number of citations that patent $n$ received by the patents belonging to category $h$ and published within $T$ years after the publication date of $n$. Formally, we define
$I_{n,h}=\frac{CIT_{n,h}}{\max_{\{i\in\text{cohort of $n$}\}}CIT_{i,h}}\,.$
Note that the number of forward citations a patent gets can differ greatly depending on the publication year and the category it comes from. This is why we decided to share the approach adopted in Squicciarini et al. (2013) and normalize this number with respect to the cohort of the patent considered. In other words, the index we use is a disaggregation of the forward citation index, say $I_{S}(n)$, in Squicciarini et al. (2013): indeed, we have
$I_{S}(n)=\frac{CIT_{n}}{\max_{\{i\in\text{cohort of $n$}\}}CIT_{i}}=\sum_{h\in{\mathcal{H}}}\tfrac{\max_{\{i\in\text{cohort of $n$}\}}CIT_{i,h}}{{\max_{\{i\in\text{cohort of $n$}\}}CIT_{i}}}I_{n,h}\,.$
Remark 2.1 (Invariance property of the index).
Note that, if we restrict our attention to (or we can observe) only the patents published in a certain category $h^{*}$, then, for each patent $n$ belonging to this category, the index value $I_{n,h^{*}}$ remains well defined and its value does not change with respect to the one computed if we consider all the published patents. Indeed, its definition is based only on the observation of the patents belonging to category $h^{*}$.
The above defined index $I_{n,h}$ takes values in $[0,1]$. We can fix a suitable threshold $\tau$ and call “success” (according to the terminology for Bernoulli processes) only the patents that have the index greater than $\tau$. More precisely, we give the following definition:
Definition 2.1 (success for a certain category).
Given the threshold $\tau$, we say that a patent $n$ is a success for (or in) category $h$ if the index $I_{n,h}>\tau$.
According to the above definition, if we denote by $X_{n,h}$ the random variable that is equal to $1$ if patent $n$ is a success for category $h$ and equal to 00 otherwise, we obtain a finite system of Bernoulli processes, that is $\{X^{(h)}=(X_{n,h})_{n}\,:\,h\in{\mathcal{H}}\}$. These processes are not independent, but they interact. This interaction is crucial to consider, as numerous studies have highlighted that –within innovation ecosystems– innovation processes do not evolve in isolation but often influence one another through knowledge spillovers, recombination, and co-dependence, shaping both innovation outcomes and strategic investment decisions across domains (Adner and Feiler 2019, Bondarev and Krysiak 2021, Kim and Magee 2017).
In the following sections, we are going to first introduce a new model of interacting Bernoulli processes, secondly, show some related theoretical results, and finally, verify that the observed behaviors in the available data match with the ones predicted by the theoretical results. We will also perform a statistical test on the interaction intensity. The methodology is presented in general so that it can be applied with different categories or indexes or also in different contexts.
3. Mathematical model and theoretical results: Interacting Bernoulli processes
In this section we will introduce and study a new model for a finite system of interacting Bernoulli processes. In the context described in the previous section, each Bernoulli process refers to a patent category and the time-steps $t$ corresponds to the sequential (with respect to the date of publication) patents $n$. However, unless otherwise specified, the model is here presented in a general framework so that it might be applied in other contexts, too.
Let $\mathcal{H}$ be a finite set with $N=\text{card}(\mathcal{H})$. We are going to define a system $\{X^{(h)}=(X_{t,h})_{t\in\mathbb{N}\setminus\{0\}}\,:\,h\in\mathcal{H}\}$ of Bernoulli processes with parameters evolving along time according to a reinforcement rule with interaction: specifically, for each $t\geq 0$, let $X_{t+1,1},\dots,X_{t+1,N}$ be $N$ Bernoulli random variables, that are conditionally independent given the past and such that
(3.1) $P_{t,h}$ $=P(X_{t+1,h}=1\,|\,\text{past until time-step $t$})$
$=\frac{\theta_{h}+\sum_{j\in{\mathcal{H}}}\gamma_{j,h}S_{t,j}}{c+t}$
where $\theta_{h}>0$, $c\geq\theta_{h}$ are parameters that tune the initial condition in the model dynamics (for details, see (B.1) in Appendix B) and $S_{t,j}=\sum_{n=1}^{t}X_{n,j}$ is the number of successes until time-step $t$ for process $j$. Moreover, the parameters $\gamma_{j,h}$ play the fundamental role in the dynamics of the system (see again (B.1)). We refer to the matrix $\Gamma=(\gamma_{j,h})_{j,h}$ as the interaction matrix and we assume the following conditions:
  • ($\mathcal{A}1$)
[…]
  • ($\mathcal{A}2$)
    $\Gamma$ is irreducible, that is the graph with the Bernoulli processes as nodes and with $\Gamma$ as the (weighted) adjacency matrix is strongly connected. (In Appendix C we will discuss the case of a reducible matrix $\Gamma$ [^7] This extension may be particularly relevant in contexts where the USPTO –or other patent offices—modify the structure of technological categories through reclassification (Lafond and Kim 2019, Chae and Gim 2019) thus changing the landscape of interaction among the latter.)
Hence, at each time-step $t$, the probability $P_{t,h}$ that we will have a success for process $h$ at time-step $t+1$ can have an increasing dependence, not only on the number of successes observed in process $h$ itself, according to a self-reinforcement principle (Pemantle 2007), but also on the number $S_{t,j}$ of successes observed in any other process $j$ until time-step $t$ (a property that we call cross-reinforcement) . The parameter $\gamma_{j,h}$ regulates this dependence (with $\gamma_{j,h}=0$ meaning the absence of direct dependence). In other words, for each pair $(j,h)$ of processes, the parameter $\gamma_{j,h}$ quantifies how much the appearance of a success in process $j$ induces a potential future success in process $h$. It is important to note that the assumption of irreducibility of $\Gamma$ entails an interaction in this sense among all the processes: process $j$ can reinforces process $h$ directly (the case when $\gamma_{j,h}>0$) or indirectly since the presence of at least one path that joint them.
In our context, given $t$ patents, the probability $P_{t,h}$ that the future patent $t+1$ will be a success in category $h$ can have an increasing dependence on the number of past successes in any category (self- and cross-category reinforcement) with the parameter $\gamma_{j,h}$ ruling the strength of this (direct) dependence and the assumed irreducibility of the matrix $\Gamma$ generates a global interaction among all the categories (according to the literature on knowledge spillovers among patents Jaffe et al. (2000), Thompson and Fox-Kean (2005)).
The following theoretical results describe the asymptotic behavior of the success-probability $P_{t,h}$ and of the number $S_{t,h}$ of successes along time, for each Bernoulli process $h$, with $h=1,\dots,N$. The rigorous proofs of all the theorems are collected in Appendix B.
Theorem 3.1.
Under assumptions ($\mathcal{A}1$) and ($\mathcal{A}2$), denote by $\gamma^{*}\in(0,1]$ the Perron-Frobenius eigenvalue of $\Gamma$. Then, for each $h\in{\mathcal{H}}$, we have
$t^{1-\gamma^{*}}P_{t,h}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}P_{\infty,h}\,,$
where $P_{\infty,h}$ is a finite strictly positive random variable. Moreover, for each $h,j\in{\mathcal{H}}$, the ratio $P_{\infty,h}/P_{\infty,j}$ of the above limit random variables is almost surely equal to the deterministic quantity $u_{h}/u_{j}>0$, where ${\mathbf{u}}=(u_{h})_{h}$ is the (unique up to a multiplicative non-zero constant) left eigenvector of $\Gamma$ associated to $\gamma^{*}$. Finally, the above convergence is also in quadratic mean.
This result implies that
  • in the case $\gamma^{*}<1$ (that is when ${\boldsymbol{1}}^{\top}\Gamma\neq{\boldsymbol{1}}^{\top}$) the probability of observing a success in process $h$ converges almost surely toward zero at the same rate $1/t^{(1-\gamma^{*})}$ for every process;
  • in the case $\gamma^{*}=1$ (that is when ${\boldsymbol{1}}^{\top}\Gamma={\boldsymbol{1}}^{\top}$) the probability of observing a success in process $h$ converges almost surely to a finite strictly positive random limit, which is the same for each process (note that in this case ${\mathbf{u}}$ is, up to a multiplicative non-zero constant, equal to the vector ${\boldsymbol{1}}$); in other words, at the steady state, the probability of having a success in a process is the same for each process of the system.
In the innovation literature, the first case is supported by empirical findings, which note that the absence of autocatalytic structures in innovation networks can lead to a decline in innovative output (Napolitano et al. 2018). As we will see in the sequel, also our data analysis provides an estimated value $\gamma^{*}<1$. The second case, instead, is well exemplified in Kim and Magee (2017) who empirically showed that cross-domain knowledge flows help maintain a globally balanced and sustained innovation ecosystem.
Moreover, we recall that, according to the Perron-Frobenius theory, the components of the vector ${\mathbf{u}}$ are all different from zero and with the same sign and, in graph theory, they correspond to the relative eigenvector centrality scores (with respect to $\Gamma^{\top}$). The eigenvector centrality is a measure of the importance of a node in a graph with respect to its out-links (if it is computed for the adjacency matrix) or its in-links (if it is computed for the transpose of the adjacency matrix). A high eigenvector centrality score means that the node points to, or respectively is pointed to by, many nodes with high scores. In the above theorems ${\mathbf{u}}$ is the vector of the relative eigenvector centrality scores with respect to $\Gamma^{\top}$ and, hence, by (3.1), a high value of $u_{h}$ means that the probability of having a success for process $h$ depends on the number of past successes observed in many processes who themselves have high scores.
The uncertainty—or volatility—of the innovation process within technological domains, with higher variance often associated with emerging or rapidly evolving fields, has been a subject of debate (Jalonen and Lehtonen 2011, Jalonen 2012, Allen 2013). Specifically, understanding the nature of uncertainty during the innovation process is particularly helpful to clarify how evolving technologies shape capability development within ecosystems, reinforcing or constraining firms’ ability to transform skills into core competencies (Leonard-Barton 1992, O’Connor 2008, Teece and Leih 2016). In this context, Theorem 3.1 offers valuable insights. Indeed, we obtain the following corollary.
Corollary 3.2.
Under assumptions ($\mathcal{A}1$) and ($\mathcal{A}2$), denote by $\gamma^{*}\in(0,1]$ the Perron-Frobenius eigenvalue of $\Gamma$ and by ${\mathbf{u}}=(u_{h})_{h}$ its corresponding (unique up to a multiplicative non-zero constant) left eigenvector. Then, for each $h\in{\mathcal{H}}$, we have
$t^{1-\gamma^{*}}Var[X_{t+1,h}]\longrightarrow|u_{h}|\,\alpha({\mathbf{u}})\,,$
where $\alpha({\mathbf{u}})>0$ (and such that $\alpha(C{\mathbf{u}})=\alpha({\mathbf{u}})/|C|$ for each constant $C\neq 0$). Moreover, for each pair $h\neq j$ of different processes, we get
$t^{2(1-\gamma^{*})}cov(X_{t+1,h},X_{t+1,j})\longrightarrow u_{h}u_{j}\sigma^{2}({\mathbf{u}})\,,$
where $\sigma^{2}({\mathbf{u}})>0$ (and such that $\sigma^{2}(C{\mathbf{u}})=\sigma^{2}({\mathbf{u}})/C^{2}$ for each constant $C\neq 0$), so that
$t^{1-\gamma^{*}}\rho(X_{t+1,h},X_{t+1,j})\longrightarrow\sqrt{u_{h}u_{j}}\,\tfrac{\sigma^{2}({\mathbf{u}})}{\alpha({\mathbf{u}})}\,.$
Hence, the ratio of the variances $Var[X_{t,h}]/Var[X_{t,j}]$ converges to the ratio $u_{h}/u_{j}$ of the relative eigenvector centrality scores (with respect to $\Gamma^{\top}$) of the two nodes (processes) $h$ and $j$. Moreover, the correlation coefficient between the observations related to any pair of different processes converges to zero at the rate $1/t^{(1-\gamma^{*})}$ when $\gamma^{*}<1$ and converges to the same strictly positive value when $\gamma^{*}=1$.
In short, a balanced innovation ecosystem ($\gamma^{*}=1$) supports ongoing, shared uncertainty and correlated success across technologies, which characterizes connected innovation ecosystems (Jacobides et al. 2024). In contrast, a not-balanced one ($\gamma^{*}<1$) leads to asymmetric, decoupled innovation dynamics over time. In other words, in well-connected systems, sustained uncertainty and interdependence across technological domains facilitate diverse learning paths and capability recombination. This interdependence fosters mutual reinforcement across domains, enabling firms to accumulate diverse, complementary capabilities. Conversely, in fragmented systems, limited spillovers may rigidify specialization patterns, potentially turning capabilities into rigidities over time.
From Theorem 3.1, we can also deduce the following result for the processes $(S_{t,h})_{t}$, with $h\in{\mathcal{H}}$, that represent the number of successes observed along time in each Bernoulli process $h$. It is important to note that, differently from the previous quantities $P_{t,h}$, that are not observable (due to the presence of the unknown model parameters), the quantities $S_{t,h}$ can be directly observed and so used for data analyses.
Theorem 3.3.
Under assumptions ($\mathcal{A}1$) and ($\mathcal{A}2$), denote by $\gamma^{*}\in(0,1]$ the Perron-Frobenius eigenvalue of $\Gamma$ and by ${\mathbf{u}}=(u_{h})_{h}$ its corresponding (unique up to a multiplicative non-zero constant) left eigenvector. Then, for each $h\in{\mathcal{H}}$, we have
$\frac{S_{t,h}}{t^{\gamma^{*}}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}S_{\infty,h}\,,$
where $S_{\infty,h}$ is a finite strictly positive random variable. Moreover, for each $h,j\in{\mathcal{H}}$, we have
$\frac{S_{t,h}}{S_{t,j}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\frac{u_{h}}{u_{j}}\quad\text{and}\quad\frac{S_{t,h}}{\sum_{j=1}^{N}S_{t,j}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\frac{u_{h}}{\sum_{j=1}^{N}u_{j}}\,.$
This result states that the number $S_{t,h}$ of successes for all the processes $h$ grows with the same Heaps’ exponent $\gamma^{*}\in(0,1]$: that is $S_{t,h}\stackrel{{\scriptstyle a.s.}}{{\sim}}S_{\infty,h}\,t^{\gamma^{*}}$. In addition, the ratio $S_{t,h}/S_{t,j}$ provides a strongly consistent estimator of the ratio $u_{h}/u_{j}$ of the relative eigenvector centrality scores (with respect to $\Gamma^{\top}$) of the two nodes (processes) $h$ and $j$ and the share of successes observed for process $h$ converges almost surely to the absolute eigenvector centrality score of $h$.
It is interesting to note that, in the innovation framework, Theorem 3.3 highlights how the long-run proportions of successful patents in each category reflect the connectedness and influence of that category within the broader innovation network. While this relationship is formally established in the present work, related insights have been suggested in the literature without formal proof. For example, Pichler et al. (2020) show that the innovation rate of a technological domain is influenced by the innovation rates of the domains it depends on. Similarly, Sampat and Ziedonis (2005), Zhang et al. (2025) emphasize the importance of the quality of a patent’s prior connections as well as their technological domain. In an effort to further deepen this latter aspect, we developed Theorem 3.4. As explained in the following, this theorem examines not only how many successful patents there are in each target category $h$, but also traces the source category $k$ to which each successful patent originally belonged. Namely, in our particular context, we can enrich the model assuming that the number $S_{t,k,h}$ of successes until time-step $t$ in category (process) $h$ coming from belonging category $k$ is of the form $S_{t,k,h}=\sum_{n=1}^{t}X_{n,h}Y_{n,k}$, where $Y_{n,k}$ takes value $1$ if the category of patent $n$ is $k$ and 00 otherwise and
  • ($\mathcal{A}3$)
    ${\mathbf{Y}}_{n}=(Y_{n,k})^{\top}_{k\in{\mathcal{H}}}$ is independent of ${\mathbf{X}}_{n}=(X_{n,h})^{\top}_{h\in{\mathcal{H}}}$ and of all the past until time-step $n-1$ with $P(Y_{n,k}=1)=\pi_{k}\in(0,1)$ (where $\sum_{k\in{\mathcal{H}}}\pi_{k}=1$).
Under these assumptions, we obtain the following result.
Theorem 3.4.
Under assumptions ($\mathcal{A}1$), ($\mathcal{A}2$) and ($\mathcal{A}3$), denote by $\gamma^{*}\in(0,1]$ the Perron-Frobenius eigenvalue of $\Gamma$. Then, for each pair $h,k\in{\mathcal{H}}$, we have
$\frac{S_{t,k,h}}{t^{\gamma^{*}}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\pi_{k}S_{\infty,h}$
and so $S_{t,k,h}/S_{t,j,h}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\pi_{k}/\pi_{j}$.
To summarize, Theorem 3.4 provides a formal representation of how the structure of innovation inputs shapes the distribution of successful outcomes across technological domains. Specifically, it shows that, asymptotically, the share of successful innovations in a target category mirrors the origin distribution of patent activity. This result offers entrepreneurs a valuable mechanism for understanding how ecosystems embed and propagate specialization trajectories, which is a critical insight, given that the transformation of skills into core capabilities depends fundamentally on how technological domains interact and reinforce one another over time (Teece 2009, Jacobides et al. 2018, Kim and Magee 2017).
3.1. Model implication: simulations with a shock
Figure 1. Linear behavior of the successes counter $S_{t,1}$ (red line) and $S_{t,2}$ (blue line) along $t$, in $\log_{10}-\log_{10}$ scale, in two different scenarios with a positive shock on process $1$. The two scenarios are distinguished by the type of interaction, i.e. by the form of the interaction matrix $\Gamma$ used in the simulations. The shock on process $1$ occurs at time $t_{shock}=10^{4}$ and the system is observed until time-step $t=10^{7}$. We can observe the sudden rise of $S_{t,1}$ (red line) caused by the shock and how the positive effect of the shock on process $1$ propagates to process $2$ (consequent progressive rise of the blue line). We can also note that the slope of the two lines and the distance between their intercepts are the same before and after the shock, because the interaction matrix $\Gamma$ does not change: indeed, the slope is equal to $\gamma^{*}$ and the distance between the intercepts corresponds to the quantity $\log_{10}(u_{1}/u_{2})=\log_{10}(u_{1})-\log_{10}(u_{2})$.
In order to highlight what is the implication of the fact that a system follows the provided model, we here show how the effect of a shock on a process propagates to the other ones by means of the interaction in their dynamics. A vivid example of this phenomenon is the surge in pharmaceutical innovation during the COVID-19 pandemic, which triggered subsequent innovations in other fields such as logistics (Dovbischuk 2022) and IT (Li et al. 2022), and also prompted a wave of open innovation practices through which inventors widely disseminated their contributions (Brem et al. 2021, Lee and Trimi 2021, Ho 2023). Accounting for innovation shocks is essential. As the COVID-19 surge demonstrated, when a shock occurs, society may be compelled to transform latent capabilities—such as vaccine development or its digital capabilities—into core capabilities (Li et al. 2022, Dovbischuk 2022, Oladapo et al. 2023). Moreover, the timing and adaptability of firms in responding to such shocks –with the consequent propagation of innovative responses via dynamic interactions– are critical to ensuring that newly required capabilities do not merely displace former core capabilities, which, in the aftermath, risk becoming core rigidities (Atanassova and Bednar 2022).
More precisely, we have considered a system with $N=2$ processes starting with the same initial composition: $\theta_{h}=1/2$ and $c_{h}=1$. We take into account two different scenarios:
  • mean-field interaction, that corresponds to a symmetric interaction matrix $\Gamma$ of the form
    (3.2) $\gamma_{j,h}=\begin{cases}\gamma^{*}\left(\iota/N+(1-\iota)\right)\quad&\mbox{if }j=h\\ \gamma^{*}\,\iota/N\quad&\mbox{if }j\neq h\end{cases},$
    with $\gamma^{*},\,\iota\in(0,1]$, where we have chosen $\gamma^{*}=0.7$ and $\iota=0.9$, so that
    $\Gamma=\begin{pmatrix}0.385&0.315\\ 0.315&0.385\end{pmatrix}\,;$
  • non-symmetric interaction matrix $\Gamma$ with $\gamma^{*}=0.685$ and ${\mathbf{u}}=(1.394,0.575)^{\top}$, i.e.
    $\Gamma=\begin{pmatrix}0.50&0.20\\ 0.45&0.20\end{pmatrix}.$
The shock occurs at time $t_{shock}=10^{4}$ and it acts as follows: at time $t_{shock}=10^{4}$, the probability of having a success for process $1$ is increased (that is, we give a positive shock to process $1$) by replacing the parameters $\theta_{1}=1/2$ and $c_{1}=1$ by the new ones $\theta_{shock,1}=1/2+10^{4}$ and $c_{shock,1}=1+10^{4}$. These new values of the parameters remain in the dynamics of process $1$ for all the future time-steps until $t=10^{7}$. The parameters for process $2$ are unchanged: $\theta_{2}=1/2$ and $c_{2}=1$ remain the same along all the time-steps. The number of successes for the two processes, that is $S_{t,1}$ and $S_{t,2}$, are observed. In Figure 1, we can see how the positive shock causes the number of successes for process $1$ to rise (red line) and how this positive effect propagates to the other process (blue line) by means of the interaction terms in their dynamics.
4. Data analysis
Our dataset consists of the GLOBAL PATSTAT database [^8] The dataset is maintained at the IMT School.. More specifically, we collected all the patents published in the period $[1980-2018]$, with their exact (full) date of publication and their CPC-1 category. Moreover, for each patent $n$, we know if it has been cited by subsequent patents (published in the considered period) and which are the citing patents.
[…]
Figure 2 provides the asymptotic behavior, in $\log_{10}-\log_{10}$ scale, of every process $S_{t,h}$, that represents the number of successes observed in category $h$ until time-step (patent) $t$. We can appreciate how the lines exhibit the same slope, which indicates that the processes have the same Heaps’ exponent. This is exactly in accordance with Theorem 3.3. The emergence of a common Heaps’ exponent across all categories supports the fact that these innovation processes are not evolving independently but are shaped by shared systemic dynamics. We provide here, therefore, an evidence for the co-evolution of innovation processes, suggesting that the success trajectory of one technological domain is influenced by –and in turn influences –others, thereby reflecting the structural interdependencies and shared constraints within the broader innovation ecosystem.
Figure 2. Linear behavior of the successes counter $S_{t,h}$ along $t$, in $\log_{10}-\log_{10}$ scale, for each category $h$. The lines are obtained by a least square interpolation using a subsample of size $10^{4}$ of the data set. We can appreciate how the lines exhibit the same slope, which has been estimated to be equal to $\widehat{\gamma}^{*}=0.689$. Indeed, the goodness of fit $R^{2}$ index obtained imposing a common slope is $0.969$ and so basically equal to the one obtained allowing the slopes to be different across categories, i.e. $0.972$. The emergence of a common Heaps’ exponent across all categories supports the fact that these innovation processes are not evolving independently but are shaped by shared systemic dynamics. Moreover, the sub-linear growth reflects diminishing returns to cumulative knowledge, increasing innovation complexity, and a bounded scalability in innovation outputs over time.
[…]
Figure 3. Plot of the process $S_{t,h}/S_{t,H}$ along time $t$, for all the categories $h\neq H$. Category $H$ is arbitrarily chosen to be the baseline category. The horizontal dashed red lines represent an estimation of the ratio $u_{h}/u_{H}$, obtained as $10^{d_{h,H}}$, where $d_{h,H}$ is the difference between the intercepts of the regression lines in Figure 2 for the categories $h$ and $H$. This long-run convergence across categories highlights a structural stabilization in innovation outcomes, despite increasing complexity and declining marginal returns to knowledge production. Moreover, note that all the estimated ratios $u_{h}/u_{H}$ are around $1$. This means that in the long run the numbers of successes in the different categories tend to coincide.
Figure 3 shows, for each category $h$, the convergence of the process $(S_{t,h}/S_{t,H})_{t}$ toward the quantity $u_{h}/u_{H}$, estimated as $10^{d_{h,H}}$ where $d_{h,H}$ is the difference between the intercepts of the regression lines in Figure 2 for the two processes related to the pair $(h,H)$ of categories. [^11] Category H in Figure 3 is arbitrarily chosen to be the baseline category. Any other category can be used as a baseline. Also this fact is in accordance with Theorem 3.3 and it means that the relative number of successes across categories stabilizes along time. This effect reflects the fact that, in the long run, the rates at which capabilities evolve into core capabilities—and core capabilities into rigidities—within a technological domain, including the cross-sectoral influences, tend to settle. In other words, the temporal stabilization of innovation successes implies two distinct mechanisms: (i) the convergence of the catalytic effect of general-purpose technologies on innovation across diverse domains (Trajtenberg et al. 1997), as reflected in the equilibrium configuration of the subset of $\Gamma$ related to such technologies; and (ii) the stabilization in the rate of recombination of existing knowledge, as captured by the behavior of $P_{t,h}$ over time. Moreover, the limit corresponds to the ratio of their respective components of the eigenvector-centrality scores ${\mathbf{u}}$ (with respect to $\Gamma^{\top}$). In Figure 3 we can note that the estimated quantities $u_{h}/u_{H}$ are all very near to $1$ (more precisely, they are all within the interval $[1,1.3]$). This means that in the long run the numbers of successes in the different categories tend to coincide, which carries significant implications for innovation policy. In the case of the green transition, for instance, it means that even if different technology areas seem to be producing similar results over time, this similarity can hide the fact that it’s actually getting harder and more complex to create those innovations (Bloom et al. 2020).
Figure 4. Linear behavior of the successes counter $S_{t,k,h}$ along $t$ in $\log_{10}-\log_{10}$ scale, for each pair $(k,h)$ of categories ($k=$sub-figure, $h=$color). The lines are obtained by a least square interpolation based on a subsample of the dataset with a slope equal to the previously estimated value $\widehat{\gamma}^{*}=0.689$. The goodness of fit $R^{2}$ index is $0.851$, which is basically the same as the one obtained allowing the slopes to be different across the pairs $(k,h)$ of categories, i.e. $0.855$. The common slope supports one again the presence of interaction across the categories. Moreover, the sub-linear growth reflects a consistent sub-linear scaling of cross-domain innovation flows, suggesting that while knowledge recombination persists, its marginal productivity diminishes over time across all category pairs.
Figure 4 provides the asymptotic behavior, in $\log_{10}-\log_{10}$ scale, of every process $S_{t,k,h}$, that represents the number of successes for category $h$ coming from category $k$ observed until time-step (patent) $t$. We can appreciate how the lines exhibit the same slope, equal to the previously estimated value $\widehat{\gamma}^{*}=0.689$ in accordance with Theorem 3.4, supporting once again the presence of interaction across the categories.
Figure 5. Plot of the process $S_{t,k,h}/S_{t,H,h}$ along time $t$, for all the pairs $(k,h)$ of categories ($k=$sub-figure, $h=$color), with $k\neq H$. Category $H$ is arbitrarily chosen to be the baseline category. The horizontal dashed red lines represent the value $10$ to the power of the differences between the intercepts of the regression lines in Figure 4 for the pairs of categories $(k,h)$ and $(H,h)$. This long-run proportionality links early patenting intensity to long-term innovation success, highlighting the path-dependent and self-reinforcing nature of technological trajectories.
Finally, each panel of Figure 5 refers to the successes observed in a given category $h$ and shows, for each coming category $k$, the convergence of the process $S_{t,k,h}/S_{t,H,h}$ toward the quantity $\pi_{k}/\pi_{H}$, estimated as 10 to the power of the difference between the intercepts of the corresponding regression lines in Figure 4.121212Category H in Figure 5 is arbitrarily chosen to be the baseline category. Any other category can be used as a baseline. Also this fact is coherent with Theorem 3.4. In the long run, therefore, the number of successful innovations in each category ends up matching how much patent activity started in each category. Hence, if a certain type of technology had a greater patent activity early on, it will also tend to have more successful innovations over time.
4.1. Statistical inference on the interaction intensity under the mean-field assumption
Assuming the system is subjected to a mean-field interaction (note that in this case we have $u_{h}/u_{j}=1$ for each $h,j\in{\mathcal{H}}$), i.e. assuming the interaction matrix $\Gamma$ to be of the form (3.2), we can perform a suitable test on the parameter $\iota$, that rules the intensity of the interaction. Indeed, since the leading eigenvalue $\gamma^{*}$ can be estimated with high accuracy by a suitable regression on the vector process ${\mathbf{S}}_{t}=(S_{t,h})_{h\in{\mathcal{H}}}^{\top}$ (see Aletti et al. (2023a) and references therein), we can assume this parameter to be known and we can make inference only on the intensity parameter $\iota$. For a two-sided test with $H_{0}:\;\iota=\iota_{0}$, where $\iota_{0}\leq 1$ and $\iota_{0}>1/2$ (the required condition in the theoretical result for having the second eigenvalue of $\Gamma$ strictly smaller than $\gamma^{*}/2$), we can use the test statistic (see Appendix B for the technical details)
(4.1) $2\Delta_{0}\frac{\big{\|}{\mathbf{S}}_{t}-\widetilde{S}_{t}{\boldsymbol{1}}\big{\|}^{2}}{\widetilde{S}_{t}}\stackrel{{\scriptstyle d}}{{\sim}}\chi^{2}(N-1)\quad\mbox{under }H_{0},$
where $\Delta_{0}=\iota_{0}-\frac{1}{2}$ and $\widetilde{S}_{t}=\tfrac{{\boldsymbol{1}}^{\top}{\mathbf{S}}_{t}}{N}=\sum_{h\in{\mathcal{H}}}S_{t,h}/N$. Since this statistic is increasing in $\iota_{0}$, it works well also for the one-sided test with $H_{0}:\;\iota\geq\iota_{0}$, with $\iota_{0}\in(1/2,1]$. Applying this one-sided test to our real dataset of patents (with the estimated value $\widehat{\gamma}^{*}=0.689$), we obtain different $p$-values, one for each tested value $\iota_{0}$, that are collected in Table 1. We can see that, if we choose the significance level $\alpha=0.05$, then the minimum value of $\iota_{0}$ at which we can reject the null hypothesis is $\iota_{0}=0.75$.
We have also computed an estimate of the parameter of interest $\iota$ maximizing the (joint) likelihood of the processes $X^{(h)}=(X_{n,h})_{n}$, $h\in{\mathcal{H}}$, i.e.
$L(\iota\,|\,x_{n,h},\,h\in{\mathcal{H}},\,n=1,\dots,n_{tot})=\prod_{n=1}^{n_{tot}}\prod_{h\in{\mathcal{H}}}P_{h,n}^{x_{h,n}}(1-P_{h,n})^{1-x_{h,n}}.$
The obtained estimate is in accordance with the results of Table 1. Indeed, the maximum-likelihood estimate of the interaction intensity $\iota$ is $\widehat{\iota}=0.643$ and from Table 1 we get that, at the level $\alpha=0.05$, we can reject the null hypothesis $H_{0}:\iota\geq 0.75$, while we do not have enough statistical evidence to reject for smaller values of $\iota_{0}$.
$\iota_{0}$ 0.55 0.6 0.65 0.7 0.75 0.8 0.85 0.9 0.95
p-value 0.893 0.562 0.273 0.113 0.042 0.015 0.005 0.002 $<0.001$
Table 1. $p$-values associated to the hypothesis test with $H_{0}:\iota\geq\iota_{0}$ under the mean-field interaction .
Our estimate $\hat{\iota}$ of the intensity parameter $\iota$ shows that on the one hand, processes are interdependent, i.e. they share knowledge, influence, or innovation spillovers and on the other, that self-dynamics still matter. This reinforces the concept of dynamic competitive advantage as introduced by Teece et al. (1997) and Giannitsis and Kager (2009). Broadly speaking, the transformation of general capabilities into core capabilities occurs through a process of smart specialization. In this process, technological capabilities must adapt to multidisciplinary stimuli originating from other technologies, translate them into their own technological language, and leverage them to explore new opportunities within their own domain.
5. Conclusions
Formalizing the mechanisms by which innovation trajectories emerge, persist, and evolve into specialized paths is essential for understanding how capabilities develop into either core competencies or core rigidities (Leonard-Barton 1992). This paper offers a formal framework to capture the dynamics of innovation ecosystems through a novel model of interacting reinforced Bernoulli processes.
Our contribution advances existing models of interacting Bernoulli processes by introducing a fully historical reinforcement mechanism with temporally decaying influence. This formulation enables a unified mathematical treatment of innovation dynamics, jointly capturing—for the first time to our knowledge—a set of empirical regularities typically addressed in isolation: sublinear growth of the number of successes, convergence of relative success rates across technological domains, and a decline in cross-category correlations. Central to our approach is the explicit modeling of both self- and cross-category reinforcement, offering a rigorous explanation for how success propagates across interdependent innovation ecosystems. The theoretical results, all proven rigorously, show that success probabilities evolve under a Perron-Frobenius regime, generating stable long-run success distributions that reflect the centrality of each domain in the interaction network. Empirical analysis of GLOBAL PATSTAT data (1980–2018) confirms these patterns, revealing a common Heaps’ exponent across categories, time-stabilizing ratios of success, and directional propagation of innovation activity. Through simulation, we also demonstrate how shocks in one domain can systematically influence others—particularly under strong coupling conditions. Moreover, we perform a statistical inference analysis under a mean-field assumption, which reveals substantial but non-maximal interaction strength across domains. Lastly, it is worth mentioning that, due to the generality of the introduced model and the related theoretical results and statistical procedures, it can be fruitfully applied to other contexts.
Beyond its theoretical contributions, the framework offers practical relevance for managers. It namely underscores the importance of early and well-targeted interventions—particularly before ecosystems fragment into highly specialized, path-dependent trajectories (David 1985, Mazzucato 2018). The model further suggests that supporting cross-domain connectivity (e.g., via interdisciplinary R&D programs or incentives for recombination across patent classes) may delay the onset of technological rigidities and foster broader capability evolution (Jacobides et al. 2018, Kim and Magee 2017). Finally, by identifying the statistical structure of innovation interactions, our model provides policymakers with a novel tool to detect when technological fields are becoming saturated—informing decisions on resource allocation and diversification strategies in national or regional innovation systems (Fleming 2001, Jaffe and Trajtenberg 2002).
Overall, this work contributes to the growing body of research aiming to formalize the dynamic interplay between innovation and capability development (see e.g. Lawson and Samson (2001), Breznik and D. Hisrich (2014)), offering a rigorous yet adaptable framework for analyzing the evolution of innovation processes in shaping innovation ecosystems.
Author contributions
All the authors contributed equally to the present work.
Acknowledgments
Giacomo Aletti is a member of the Italian Group “Gruppo Nazionale per il Calcolo Scientifico” of the Italian Institute “Istituto Nazionale di Alta Matematica”. Irene Crimaldi is a member of the Italian Group “Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni” of the Italian Institute “Istituto Nazionale di Alta Matematica”.
Fundings
Irene Crimaldi and Federico Nutarelli are partially supported by the European Union - NextGenerationEU through the Italian Ministry of University and Research under the National Recovery and Resilience Plan (PNRR) - M4C2 - Investment 1.3, title [Public Sector Indicators for Sustainability and Wellbeing (PUWELL)] - Program [Growing Resilient, INclusive and Sustainable (GRINS)] - PE18 - CUP J33C22002910001.
This research was supported by the “Resilienza Economica e Digitale” (RED) project
(CUP D67G23000060001) funded by the Italian Ministry of University and Research (MUR) as “Department of Excellence” (Dipartimenti di Eccellenza 2023-2027, Ministerial Decree no. 230/2022)
Modeling Innovation Ecosystem Dynamics through Interacting Reinforced Bernoulli Processes. Supplementary material
Appendix A Definitions of novelty in the literature
For a patent to be defined as such, it must be “novel” at the time of its filing (Mueller 2024). In other words, a patent is novel by definition, and the concept of novelty is widely discussed in the literature (Abbas et al. 2014). Is novelty related to how impactful a patent is? Or how influential it becomes? Or perhaps how unexpected its release was?
Source Novelty Index Quick explanation Based on
Squicciarini et al. (2013) $\text{Originality}_{p}=1-\sum_{j}^{n_{p}}s_{pj}^{2}$ $s_{pj}$ is the percentage of citations made by patent p to patent class $j$ out of the $n_{p}$ IPC 4-digit (or 7-digit) patent codes contained in the patents cited by patent $p$. Citation measures are built on EPO patents and account for patent equivalents. Citations
Squicciarini et al. (2013) $\text{Radicalness}_{p}=\sum_{j}^{n_{p}}\frac{\text{CT}_{j}}{n_{p}}\,;\,\text{IPC}_{pj}\neq\text{IPC}_{p}$ $CT_{j}$ denotes the count of IPC-4 digit codes $CPC_{pj}$ of patent $j$ cited in patent $p$ that is not allocated to patent $p$, out of $n$ CPC classes in the backward citations counted at the most disaggregated level available (up to the 5th hierarchical level). The higher the ratio, the more diversified the array of technologies on which the patent relies upon. Citations Tech.Knowledge
Own elaboration based on intuitions in Hall et al. (2000) and Lanjouw and Schankerman (2004) $\text{ANPCI}_{i,t}=\frac{C_{i,t}}{\lambda_{t}\cdot\left(\frac{1}{n_{t}}\sum_{j=1}^{n_{t}}C_{j,t}\right)}$ where: - $C_{i,t}$ is the number of citations received by patent $i$ in year $t$, - $n_{t}$ is the number of patents filed in year $t$, - $\frac{1}{n_{t}}\sum_{j=1}^{n_{t}}C_{j,t}$ is the average number of citations per patent in year $t$, -$\lambda_{t}$ is a correction factor to account for citation inflation, which adjusts the average citation counts for inter-year comparability. Is the average number of citations per patent across all fields in year $t$, divided by a baseline year (e.g., $t_{0}$) Intra-year comparison: index is normalized by the average citations within the same cohort (year and field). Inter-year comparison: adjusting for citation inflation $\lambda_{t}$ with , we should mitigate the bias of comparing older patents with more citations to newer ones that haven’t had as much time to accumulate citations. Citations
Trajtenberg et al. (1997) (Modified better version) $G_{X}=1-\sum_{j=1}^{M_{i}}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{T_{ji}^{n}}{T_{i}^{n}}\right)^{2}$ $X$ is the focal patent with $Y_{i}$ patents citing the focal patent $X$, with $i=1,\cdot,N$. $T_{i}^{n}$ is the total number of IPC $n$-digit classes in $y_{i}$ $T_{ji}^{n}$ is the total number of IPC $n$-digit classes in the $j^{th}$ IPC4 digit class in $y_{i}$ and $j=1\dots M_{i}$ is the cardinal of all IPC4-digit classes in $y_{i}$ Citations Tech.Knowledge
Various Patent quality: Composite indices Typically based on patent citations, claims, patent renewals and patent family size. Usually different compositions as follows: [i]Patent quality index 4 – 4 components: number of forward citations (up to 5 years after publication); patent family size; number of claims; and the patent generality index. Only granted patents are covered by the index. [ii] Patent quality index 4b – 4 components, bis: number of forward citations (up to 5 years after publication); patent family size; corrected claims; and the patent generality index. Only granted patents are covered by the index. [iii] Patent quality index 6 – 6 components: covers the same components as above, plus the number of backward citations and the grant lag index. Entropy based approaches: Defines a weighting scheme ($W$) for all indicators of a patent (citations, claims…) based on entropy. With $M$ patents, it then constructs $M\cdot W$. A set of negative (patents with the lowest weighted scores across the majority of indicators) and positive patents is then identified and divided; a similarity across good and negative patents is then computed. The following steps are then performed: Maximum Similarity: For each patent in $M^{R}$, the maximum similarity to any patent in $M^{N}$ is determined. If the maximum similarity $S(M_{i}^{R},M^{N})$ exceeds a predefined threshold $\tau$, that remaining patent is marked as negative. The process repeats iteratively. Newly marked negative patents are moved to $M^{N}$, and the model re-evaluates the remaining patents in $M^{R}$. The iteration stops when the number of remaining patents $M^{R}$ is less than a threshold $\theta$, or no new patents are marked as negative. The final output is a small set of patents in $M^{R}$ that are not marked as negative, indicating their higher potential for technological innovation based on distinctiveness and low similarity to the negative patents. Various
Table 2. This Table lists the various novelty indices brought by the literature. As noted in Lanjouw and Schankerman (2004), however, no single indicator can fully capture the quality or value of a patent and multiple indicators (citations, claims, oppositions, family size) provide a more comprehensive and nuanced understanding of both the technological importance and the economic potential of innovations.
Thus, the papers that address “patent novelty” adopt their own definitions of the term. Consequently, if we attempt to classify the main streams of literature on patent novelty, we may identify the following:
  • i)
    The older and more established literature is the one that defines novelty basing on the forward and backward citations (Trajtenberg et al. 1997, Squicciarini et al. 2013). Moreover, in the document about PATSTAT database we can read: “The number of citing patent documents can be an indicator of the importance of the patent. A frequently cited patent can be an indication of a core technology”. This literature have been recently criticized. In Abbas et al. (2014), for instance, the authors primarily discuss patent novelty in the context of citation analysis, classification systems, and keywords. They highlight how backward citations (patents cited by a new patent) are commonly used to infer novelty—specifically, fewer backward citations can signal higher novelty because the patent draws less from existing technologies. However, they criticize citation-based methods, noting that they don’t always capture the full semantic or functional novelty of a patent.
  • ii)
    In Fleming (2001) novelty is primarily defined through the recombination of existing knowledge. The core idea is that innovations arise by combining different components (Fleming uses U.S. patent classification codes as a proxy for these components by examining how frequently different combinations of these classifications appear together in patents, Fleming can assess the novelty of the recombination [^13] Specifically, he looks at how often certain technological combinations have been used before. If a patent combines classifications that rarely appear together or haven’t been combined before, it’s considered more novel. This method helps to quantify novelty based on how “new” or “unusual” the combination of technological elements is, rather than just looking at the number of citations or backward links. Fleming also explores the uncertainty associated with novelty, showing that the more novel the combination (i.e., the less frequently those components have been combined in the past), the greater the uncertainty in terms of the patent’s success.) in new ways. A patent is considered novel when it brings together elements (or knowledge) that haven’t been combined before, thus creating something new. Fleming emphasizes that recombinant novelty can vary in terms of uncertainty and outcomes: some combinations lead to breakthroughs, while others fail to generate impactful results. Fleming also makes a distinction between simple recombination, which involves putting together elements that are closely related or have been combined before, and more radical recombination, where the components are from different, often unrelated, technological fields. The latter leads to higher novelty and a greater chance of disruption but also comes with more uncertainty in terms of success. Thus, novelty in this context is seen as the creation of new combinations of existing components, with the degree of novelty depending on how distant or different these components are from one another.
  • iii)
    In the early $2000$s also another stream of literature developed following the influential publication of Aghion et al. (2005) according to which novelties were characterized by a certain degree of creative destruction. This idea has been incorporated in some studies on patents the most insightful of which is maybe Autor et al. (2020). The idea, though not explicitly stated, is that a novel patent is one that “destroys” similar patents (e.g. reducing the sales of products sold with patents in the same sector). Though this idea is very appealing it cannot be easily implemented.
  • iv)
    The most recent stream is the one of Natural Language Programming (NPL) and text analyses of patents (Abbas et al. 2014, Gerken and Moehrle 2012). In essence in Gerken and Moehrle (2012) the process of calculating patent novelty compares the text of a new patent with all previously filed patents. The similarity between the new patent and each prior one is measured, and the highest similarity score (the patent’s “oldness”) is subtracted from $1$ to determine the novelty score. Essentially, the more distinct a new patent is from earlier patents, the higher its novelty. This approach emphasizes the unique content of a patent by minimizing similarities with existing technologies. In Abbas et al. (2014), semantic analysis is mentioned as a developing method, with an emphasis on how linguistic analysis (e.g., analyzing keywords, abstracts) is becoming crucial in identifying novel patents. But, their definition of novelty primarily revolves around citation gaps [^14] If a patent doesn’t reference many previous patents (or none at all), it suggests that the new patent is introducing ideas not heavily reliant on earlier work. This is seen as a sign of novelty because it indicates that the patent isn’t just building on existing knowledge, but possibly presenting something new. and technological distance [^15] This refers to how different a patent is from those in similar fields. If a patent belongs to a new or less crowded technological class (based on how patents are categorized), it suggests that the patent is more novel because it explores a technological area that hasn’t been fully developed yet. (e.g., patents in new technological classes). More recently Kelly et al. (2021) defined, again, patent novelty by focusing on textual similarity. They measure novelty by comparing the text content of patents to both previous and future patents. A patent is considered novel if it is distinct from earlier patents but also relevant to future innovations. Their approach uses textual analysis to identify patents that contribute significantly to technological progress, and these are often predictive of future citations and market value.
Table 2 summarizes the above indices. Note that different indices may capture different aspects of a patent. For example, the forward citations of a patent $n$ measure its impact on subsequent “production” or innovation, whereas the composition of patent $n$ across categories reflects its degree of “interdisciplinarity”, which can be interpreted as a form of “originality”. Therefore, it is not necessarily the case that all indices are positively correlated with one another.
In the main text we adopt a category-version of the Squicciarini et al. (2013) index as our primary measure of patent novelty due to its widespread recognition and methodological robustness within the economics of innovation literature (Higham et al. 2021, Dehghani et al. 2023). This index, endorsed and implemented by the OECD in its patent quality indicators framework, captures the diversity of technological fields cited by a patent-interpreted as a proxy for knowledge recombination and inventive originality. It has also been applied across a range of recent empirical studies (Dehghani et al. 2023, Gao and Lazarova 2023, Angori et al. 2024) reflecting its growing standardization and comparability across datasets and time. Compared to newer text-based or machine-learning approaches, the Squicciarini et al. (2013)’s index offers a better replicability, and ease of modification in a cross-category setting as ours. Notice, however that our setting is general enough to allow for other measures of patent novelty as well.
Appendix B Proofs of the main results and technical details for the statistical test
We here provide the rigorous proofs of the main results and the technical details that explain the performed statistical test. These proofs are based on the general theory we develop in Subsection B.1 and which has the merit to provide a general mathematical framework that can be applied also in other settings.
Proof.
Proof of Theorem 3.1 For each time step $t$, set ${{\mathbf{X}}}_{t}=(X_{t,1},\dots,X_{t,N})^{\top}$, ${{\mathbf{P}}}_{t}=(P_{t,1},\dots,P_{t,N})^{\top}$ and ${{\boldsymbol{\theta}}}=(\theta_{1},\dots,\theta_{N})^{\top}$. Hence, the vectorial dynamics for the random vector ${{\mathbf{P}}}_{t}$ is, for $t\geq 0$,
[…]
${{\mathbf{P}}}_{t+1}$ $=\Big{(}1-\frac{1}{t+1}\Big{)}{{\mathbf{P}}}_{t}+\frac{1}{t+1}\Gamma^{\top}{{\mathbf{P}}}_{t+1}+O(1/t^{2}){\boldsymbol{1}}$
$={{\mathbf{P}}}_{t}-\frac{1}{t+1}(I-\Gamma^{\top}){{\mathbf{P}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}^{\prime}_{t+1}+O(1/t^{2}){\boldsymbol{1}},$
where $\Delta{{\mathbf{M}}}^{\prime}_{t+1}={{\mathbf{X}}}_{t+1}-{{\mathbf{P}}}_{t}$.
Now, we fix $x>0$ and set
(B.2) $\zeta_{0}(x)=1,\qquad\zeta_{t}(x)=\frac{\Gamma(t+x)}{\Gamma(t)}\sim t^{x}\uparrow+\infty\,.$
More precisely, from (Gouet 1993, Lemma 4.1) we have
$\zeta_{t}(x)=t^{x}+O(t^{x-1})\quad\mbox{and}\quad\frac{1}{\zeta_{t}(x)}=\frac{1}{t^{x}}+O(1/t^{x+1}),$
and so
(B.3) $\frac{1}{\zeta_{t+1}(x)\zeta_{t+1}(1-x)}=$
$\Big{(}\frac{1}{(t+1)^{x}}+O(1/t^{x+1})\Big{)}\Big{(}\frac{1}{(t+1)^{1-x}}+O(1/t^{2-x})\Big{)}=$
$\frac{1}{t+1}+O(1/t^{2}).$
Hence, multiplying (B.1) by $\zeta_{t+1}(1-\gamma^{*})$ and using the relation
$\frac{\zeta_{t+1}(x)}{\zeta_{t}(x)}=\frac{\Gamma(t+x+1)}{\Gamma(t+1)}\frac{\Gamma(t)}{\Gamma(t+x)}=1+\frac{x}{t}=1+\frac{x}{t+1}+O(1/t^{2}),$
with $x=1-\gamma^{*}$, we get the following dynamics for ${{\mathbf{A}}}_{t}=\zeta_{t}(1-\gamma^{*}){{\mathbf{P}}}_{t}$, where we set $\Delta{{\mathbf{M}}}_{t+1}=\zeta_{t+1}(1-\gamma^{*})\Delta{{\mathbf{M}}}^{\prime}_{t+1}$:
${{\mathbf{A}}}_{t+1}$ $=\Big{(}1-\frac{1}{t+1}(I-\Gamma^{\top})\Big{)}\frac{\zeta_{t+1}(1-\gamma^{*})}{\zeta_{t}(1-\gamma^{*})}{{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\Big{(}\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\Big{)}{\boldsymbol{1}}$
$=\Big{(}1-\frac{1}{t+1}(I-\Gamma^{\top})\Big{)}\Big{(}1+\frac{1-\gamma^{*}}{t+1}+O(1/t^{2})\Big{)}{{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\Big{(}\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\Big{)}{\boldsymbol{1}}$
$={{\mathbf{A}}}_{t}-\frac{1}{t+1}(\gamma^{*}I-\Gamma^{\top}){{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\Big{(}\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\Big{)}{\boldsymbol{1}}\,,$
with $O(\zeta_{t}(1-\gamma^{*})/t^{2})=O(1/t^{1+\gamma^{*}})$.
Now, we are going to apply Theorem B.2 of Section B.1 with $\Phi=\Gamma$, $\phi^{*}=\gamma^{*}$ and ${\mathcal{F}}=({\mathcal{F}}_{t})_{t}$ the natural filtration associated to the model, i.e. ${\mathcal{F}}_{t}=\sigma(X_{n,h}:\,n\leq t,\,h=1,\dots,N)$. To this purpose, we choose ${\mathbf{u}}$ and ${\mathbf{v}}$ as the left and the right eigenvectors of $\Gamma$ associated to $\gamma^{*}$, with strictly positive components and such that ${\mathbf{v}}^{\top}{\boldsymbol{1}}=1$ and ${\mathbf{v}}^{\top}{\mathbf{u}}=1$. (Recall that $\gamma^{*}$ is simple and it is possible to choose the components of these vectors all strictly positive because of the Frobenius-Perron theory.) We set $\widetilde{A}_{t}={\mathbf{v}}^{\top}{\mathbf{A}}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{v}}^{\top}{\mathbf{P}}_{t}$. First of all, we observe that assumptions (i) and (ii) of Theorem B.2 are satisfied because of ($\mathcal{A}1$) and ($\mathcal{A}2$), that imply $\gamma^{*}\in(0,1]$. Moreover, also assumption (iii) is verified: indeed, we have $\sup_{t}E[\widetilde{A}_{t}]<+\infty$ as $\widetilde{A}_{t}$ is non-negative (see Remark B.1). Finally, we have
$W_{t}=\sum_{j=1}^{N}E[(\Delta M_{t+1,j})^{2}\,|\,\mathcal{F}_{t}]=\zeta_{t+1}(1-\gamma^{*})^{2}\sum_{j=1}^{N}P_{j,t}(1-P_{j,t})\leq\zeta_{t+1}(1-\gamma^{*})^{2}\sum_{j=1}^{N}P_{j,t}.$
Then, denoting by $v_{\min}>0$ the minimum element of ${\mathbf{v}}$, we obtain
$\zeta_{t+1}(1-\gamma^{*})^{2}\sum_{j=1}^{N}P_{j,t}\leq\zeta_{t+1}(1-\gamma^{*})^{2}{\mathbf{v}}^{\top}{\mathbf{P}}_{t}/v_{\min}\leq 2\zeta_{t+1}(1-\gamma^{*})\widetilde{A}_{t}/v_{\min}.$
and so
$W_{t}\leq\frac{1}{\gamma^{*}v_{\min}}\zeta_{t+1}(1-\gamma^{*})\widetilde{A}_{t}\,.$
Therefore, recalling that $\sup_{t}E[\widetilde{A}_{t}]<+\infty$ as told before, we get
$\sum_{t}\frac{1}{(t+1)^{2}}E[W_{t}]\leq\frac{1}{\gamma^{*}v_{\min}}\sup_{t}E[\widetilde{A}_{t}]\sum_{t}\frac{\zeta_{t+1}(1-\gamma^{*})}{(t+1)^{2}}<+\infty.$
Hence, also assumption (iv) in Theorem B.2 is satisfied and we get
$t^{1-\gamma^{*}}{\mathbf{P}}_{t}\stackrel{{\scriptstyle a.s./L^{2}}}{{\sim}}\zeta_{t}(1-\gamma^{*}){\mathbf{P}}_{t}={\mathbf{A}}_{t}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}{\mathbf{P}}_{\infty}=\widetilde{P}_{\infty}{\mathbf{u}}\,,$
where $\widetilde{P}_{\infty}=\widetilde{A}_{\infty}$ is a square-integrable non-negative random variable, which is the almost sure limit of the above defined $\widetilde{A}_{t}$, or equivalently (by (B.2)), of $t^{1-\gamma^{*}}{\mathbf{v}}^{\top}{\mathbf{P}}_{t}$
It remains to prove that $P(\widetilde{P}_{\infty}>0)=1$. To this purpose, we use (Aletti et al. 2023a, Theorem S1.3). Indeed, setting $\widetilde{P}_{t}={\mathbf{v}}^{\top}{\mathbf{P}}_{t}$, we have
(B.4) $\widetilde{P}_{0}$ $={\mathbf{v}}^{\top}{\mathbf{P}}_{0}=\frac{1}{c}{\mathbf{v}}^{\top}\boldsymbol{\theta}>0$
$\widetilde{P}_{t+1}$ $=\Big{(}1-\frac{1}{t+1}\Big{)}\widetilde{P}_{t}+\frac{\gamma^{*}}{t+1}\widetilde{X}_{t+1}+O(1/t^{2}),\quad t\geq 0,$
with $\widetilde{X}_{t+1}={\mathbf{v}}^{\top}{\mathbf{X}}_{t+1}$. Hence, if we define the stochastic process $\mathcal{V}=(\mathcal{V}_{t})_{t\geq 0}$, taking values in the interval $[0,1]$, as
$\mathcal{V}_{0}$ $=\widetilde{P}_{0}>0,\qquad\qquad\mathcal{V}_{t+1}=\Big{(}1-\frac{1}{t+2}\Big{)}\mathcal{V}_{t}+\frac{1}{t+2}Y_{t+1},\quad t\geq 0,$
where $Y_{t+1}=\gamma^{*}\widetilde{X}_{t+1}$ (that takes values in $[0,1]$, since $0<\gamma^{*}\leq 1$, $X_{t+1,j}\in\{0,1\}$ and $\mathbf{v}^{\top}\mathbf{1}=1$), then by the technical result in Subsection B.2 (applied to $\mathcal{W}_{t}=\widetilde{P}_{t}$ with $\beta=1$) we have
$|\widetilde{P}_{t}-\mathcal{V}_{t}|=O(\ln(t)/t)\to 0$
and also
$t^{1-\gamma^{*}}|\widetilde{P}_{t}-\mathcal{V}_{t}|=O(t^{1-\gamma^{*}}\ln(t)/t)=O(\ln(t)/t^{\gamma^{*}})\to 0\,.$
Hence, from (Aletti et al. 2023a, Theorem S1.3) applied to ${\mathcal{V}}=(\mathcal{V}_{t})$ with $\delta=\gamma^{*}$ (note that $E[Y_{t+1}|\mathcal{F}_{t}]=\gamma^{*}\widetilde{P}_{t}=\gamma^{*}\mathcal{V}_{t}+O(\ln(t)/t)$, we get that $t^{1-\gamma^{*}}\mathcal{V}_{t}$ converges almost surely to a strictly positive finite random variable. This random variable is obviously also the almost sure limit of $t^{1-\gamma^{*}}\widetilde{P}_{t}$ and so we can conclude that $P(\widetilde{P}_{\infty}>0)=1$. ∎∎
Proof.
Proof of Corollary 3.2 In the previous proof, we have proven that the limit random vector of $t^{1-\gamma^{*}}{\mathbf{P}}_{t}$ is ${\mathbf{P}}_{\infty}=\widetilde{P}_{\infty}{\mathbf{u}}$, where here $\widetilde{P}_{\infty}$ refers to a precise choice for the vector ${\mathbf{u}}$, that is $\widetilde{P}_{\infty}=\widetilde{P}_{\infty}({\mathbf{u}})$. If we choose a different (left) eigenvector ${\mathbf{u}}^{\prime}$ of $\Gamma$ associated to $\gamma^{*}$, then we necessarily have have ${\mathbf{u}}^{\prime}=C{\mathbf{u}}$ with $C\neq 0$ and so we can write the limit random vector as ${\mathbf{P}}_{\infty}=\widetilde{P}_{\infty}({\mathbf{u}}^{\prime}){\mathbf{u}}^{\prime}$ with $\widetilde{P}_{\infty}({\mathbf{u}}^{\prime})=\widetilde{P}_{\infty}({\mathbf{u}})/C$. Summing up, for any choice of the (left) eigenvector ${\mathbf{u}}$, the limit random vector can be written as ${\mathbf{P}}_{\infty}=\widetilde{P}_{\infty}({\mathbf{u}}){\mathbf{u}}$ with a suitable square-integrable random variable $\widetilde{P}_{\infty}({\mathbf{u}})$ such that $P(\widetilde{P}_{\infty}({\mathbf{u}})\neq 0)=1$ and $\widetilde{P}_{\infty}(C{\mathbf{u}})=\widetilde{P}_{\infty}({\mathbf{u}})/C$. In other words, each random variable $P_{\infty,h}$ can be factorized in the product of a deterministic term specific for each $h$, i.e. $u_{h}$, and a common random term, i.e. $\widetilde{P}_{\infty}({\mathbf{u}})$.
Since $X_{t+1,h}\in\{0,1\}$ with $E[X_{t+1,h}\,|\,{\mathcal{F}}_{t}]=P_{t,h}$ and the convergence of $t^{1-\gamma^{*}}{\mathbf{P}}_{t}$ toward ${\mathbf{P}}_{\infty}$ is also in quadratic mean, we obtain
$t^{1-\gamma^{*}}Var[X_{t+1,h}]=t^{1-\gamma^{*}}E[P_{t,h}(1-P_{t,h})]\to\begin{cases}E[P_{\infty,h}]=u_{h}E[\widetilde{P}_{\infty}({\mathbf{u}})]\;&\text{for}\gamma^{*}<1\\ E[P_{\infty,h}(1-P_{\infty,h})]=u_{h}E\left[\widetilde{P}_{\infty}({\mathbf{u}})(1-u_{h}\widetilde{P}_{\infty}({\mathbf{u}}))\right]\;&\text{for }\gamma^{*}=1\,.\end{cases}$
Therefore, since the sign of the components of ${\mathbf{u}}$ (which is the same for all of them) necessarily coincides with the one of $\widetilde{P}_{\infty}({\mathbf{u}})$, in order to obtain the first limit relation in Corollary 3.2, we can set $\alpha({\mathbf{u}})$ equal to $|E[\widetilde{P}_{\infty}(u)]|$ when $\gamma^{*}<1$ and equal to $|E\left[\widetilde{P}_{\infty}({\mathbf{u}})(1-u_{h}\widetilde{P}_{\infty}({\mathbf{u}}))\right]|$ when $\gamma^{*}=1$, so that the two possible cases in the above formula can be summarized as $|u_{h}|\alpha({\mathbf{u}})$. Moreover, for each pair $h\neq j$, since $X_{t+1,h}$ and $X_{t+1,j}$ are conditionally independent given ${\mathcal{F}}_{t}$, we get
$t^{2(1-\gamma^{*})}cov(X_{t+1,h},X_{t+1,j})=t^{2(1-\gamma^{*})}cov(P_{t,h},P_{t,j})\longrightarrow cov(P_{\infty,h},P_{\infty,j})=u_{h}u_{j}Var[\widetilde{P}_{\infty}({\mathbf{u}})]\,.$
Hence, in order to obtain the second desired relation, we can set $\sigma^{2}({\mathbf{u}})=Var[\widetilde{P}_{\infty}({\mathbf{u}})]$. Finally, as a consequence, for the correlation coefficients, we have
$t^{1-\gamma^{*}}\rho(X_{t+1,h},X_{t+1,j})=\frac{t^{2(1-\gamma^{*})}cov(X_{t+1,h},X_{t+1,j})}{\sqrt{t^{(1-\gamma^{*})}Var[X_{t+1,h}]}\sqrt{t^{(1-\gamma^{*})}Var[X_{t+1,j}]}}\longrightarrow\sqrt{u_{h}u_{j}}\tfrac{\sigma^{2}({\mathbf{u}})}{\alpha({\mathbf{u}})}\,.$
Proof.
Proof of Theorem 3.3 From Theorem 3.1, using the same notation adopted in its proof, we get
$S_{t,h}=\sum_{n=1}^{t}X_{n,h}\qquad\mbox{with }E[X_{t+1,h}\,|\,{\mathcal{F}}_{t}]=P_{t,h}\stackrel{{\scriptstyle a.s.}}{{\sim}}\frac{P_{\infty,h}}{t^{1-\gamma^{*}}}$
and so, by (Williams 1991, sec. 12.15), we get
(B.5) $S_{t,h}\stackrel{{\scriptstyle a.s.}}{{\sim}}S_{\infty,h}\,t^{\gamma^{*}}~{}\mbox{ with }S_{\infty,h}=\frac{P_{\infty,h}}{\gamma^{*}}\,.$
(We can also note that, similarly, since the convergence in Theorem 3.1 is also in mean, we also have $E[S_{t,h}]=\sum_{n=0}^{t}E[X_{n,h}]=\sum_{n=0}^{t}E[P_{n,h}]\sim t^{\gamma^{*}}E[S_{\infty,h}]$.) As a consequence of (B.5) and the fact that $P_{\infty,h}/P_{\infty,j}=u_{h}/u_{j}$, we obtain
$\frac{S_{t,h}}{S_{t,j}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\frac{S_{\infty,h}}{S_{\infty,j}}=\frac{u_{h}}{u_{j}}\,.$
Proof.
Proof of Theorem 3.4 From Theorem 3.1 and Theorem 3.3, using the same notation adopted in their proofs, we have by ($\mathcal{A}3$)
$E[X_{t+1,h}Y_{t+1,k}\,|\,{\mathcal{F}}_{t}]$ $=E[Y_{t+1,k}]E[X_{t+1,h}\,|\,{\mathcal{F}}_{t}]=\pi_{k}P_{t,h}$
$\stackrel{{\scriptstyle a.s.}}{{\sim}}t^{-(1-\gamma^{*})}\pi_{k}P_{\infty,h}=t^{-(1-\gamma^{*})}\pi_{k}\gamma^{*}S_{\infty,h}$
and so it is enough to apply (Williams 1991, sec. 12.15) in order to obtain $S_{t,k,h}=\sum_{n=1}^{t}X_{n,h}Y_{n,k}\stackrel{{\scriptstyle a.s.}}{{\sim}}\pi_{k}S_{\infty,h}\,t^{\gamma^{*}}$. ∎∎
Technical details for the statistical inference (Subsec. 4.1)
We here use the same choice of the eigenvectors ${\mathbf{v}}$ and ${\mathbf{u}}$ and the same notation adopted in the proof of Theorem 3.1. We recall that the dynamics of the above defined process ${\mathbf{A}}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{P}}_{t}$ with $\zeta_{t}(\cdot)$ defined in (B.2), is
${\mathbf{A}}_{0}$ $={\boldsymbol{\theta}}/c$
${{\mathbf{A}}}_{t+1}$ $={{\mathbf{A}}}_{t}-\frac{1}{t+1}(\gamma^{*}I-\Gamma^{\top}){{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\left(\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\right){\boldsymbol{1}}\,,$
where $O(\zeta_{t}(1-\gamma^{*})/t^{2})=O(1/t^{1+\gamma^{*}})$ and $\Delta{{\mathbf{M}}}_{t+1}=\zeta_{t+1}(1-\gamma^{*})\Delta{{\mathbf{M}}}^{\prime}_{t+1}$ with $\Delta{{\mathbf{M}}}^{\prime}_{t+1}={{\mathbf{X}}}_{t+1}-{{\mathbf{P}}}_{t}$. Moreover, setting ${\mathbf{S}}_{t}=(S_{t,1},\dots,S_{t,N})$ and ${\mathbf{B}}_{t}=\frac{1}{\zeta_{t}(\gamma^{*})}{\mathbf{S}}_{t}$, we find the following vectorial dynamics:
${\mathbf{B}}_{0}$ $={\boldsymbol{0}}$
${\mathbf{B}}_{t+1}$ $=\frac{\zeta_{t}(\gamma^{*})}{\zeta_{t+1}(\gamma^{*})}{\mathbf{B}}_{t}+\frac{1}{\zeta_{t+1}(\gamma^{*})}{\mathbf{X}}_{t+1}=\Big{(}1-\frac{\gamma^{*}}{t+1}\Big{)}{\mathbf{B}}_{t}+\frac{1}{\zeta_{t+1}(\gamma^{*})}\Delta{\mathbf{M}}^{\prime}_{t+1}+\frac{1}{\zeta_{t+1}(\gamma^{*})}{\mathbf{P}}_{t}$
$=\Big{(}1-\frac{\gamma^{*}}{t+1}\Big{)}{\mathbf{B}}_{t}+\frac{1}{\zeta_{t+1}(\gamma^{*})\zeta_{t+1}(1-\gamma^{*})}\Delta{\mathbf{M}}_{t+1}+\frac{1}{\zeta_{t+1}(\gamma^{*})\zeta_{t+1}(1-\gamma^{*})}\frac{\zeta_{t+1}(1-\gamma^{*})}{\zeta_{t}(1-\gamma^{*})}{\mathbf{A}}_{t}\,.$
Using (B.3) and the relation $\zeta_{t+1}(x)/\zeta_{t}(x)=1+O(1/t)$, we obtain
(B.6) ${\mathbf{B}}_{t+1}$ $={\mathbf{B}}_{t}-\frac{1}{t+1}\left(\gamma^{*}{\mathbf{B}}_{t}-{\mathbf{A}}_{t}\right)$
$\qquad+\frac{1}{t+1}\Delta{\mathbf{M}}_{t+1}+O(\zeta_{t}(1-\gamma^{*})/t^{2}){\boldsymbol{1}}\,,$
where again $O(\zeta_{t}(1-\gamma^{*})/t^{2})=O(1/t^{1+\gamma^{*}})$. Finally, we observe that, by (B.2), we have
$t^{-(1-\gamma^{*})}E[\Delta{\mathbf{M}}_{t+1}\Delta{\mathbf{M}}_{t+1}^{\top}\,|\mathcal{F}_{t}]$ $\stackrel{{\scriptstyle a.s.}}{{\sim}}t^{1-\gamma^{*}}diag\left(P_{t,1}(1-P_{1,t}),\dots,P_{t,N}(1-P_{t,N})\right)$
$\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\begin{cases}diag({\mathbf{P}}_{\infty})\quad&\mbox{for }\gamma^{*}<1\\ diag(P_{\infty,1}(1-P_{\infty,1}),\dots,P_{\infty,N}(1-P_{\infty,N}))\quad&\mbox{for }\gamma^{*}=1\;(\mbox{and so }{\mathbf{u}}={\boldsymbol{1}})\end{cases}$
=$\quad=\begin{cases}\widetilde{P}_{\infty}diag({\mathbf{u}})\quad&\mbox{for }\gamma^{*}<1\\ \widetilde{P}_{\infty}(1-\widetilde{P}_{\infty})I\quad&\mbox{for }\gamma^{*}=1\;(\mbox{and so }{\mathbf{u}}={\boldsymbol{1}})\end{cases}$
and we recall that, as proven before, $\widetilde{A}_{t}={\mathbf{v}}^{\top}{\mathbf{A}}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{v}}^{\top}{\mathbf{P}}_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}\widetilde{P}_{\infty}$. Therefore, the pair $({\mathbf{A}}_{t},{\mathbf{B}}_{t})_{t}$ satisfies the dynamics and the conditions required in the general central limit theorem proven in (Aletti et al. 2025, Appendix A), provided we assume ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}<1/2$, where $\gamma^{*}_{2}$ is an eigenvalue of $\Gamma$ different from $\gamma^{*}$ with highest real part, that is $\gamma^{*}_{2}\in Sp(\Gamma)\setminus\{\gamma^{*}\}$ with ${\mathcal{R}e}(\gamma^{*}_{2})=\max\{{\mathcal{R}e}(\gamma):\gamma\in Sp(\Gamma)\setminus\{\gamma^{*}\}\}$. Hence, we can apply the statistical tools based on that result and described in (Aletti et al. 2025, Appendix B and C). In particular, in the mean-field case we have ${\mathbf{v}}=N^{-1}{\boldsymbol{1}}$, ${\mathbf{u}}={\boldsymbol{1}}$ and $\gamma^{*}_{2}=\gamma^{*}(1-\iota)$ so that, when $\iota>1/2$, we have
$(2\iota-1)t^{\gamma^{*}}\frac{\|{\mathbf{B}}_{t}-{\mathbf{v}}^{\top}{\mathbf{B}}_{t}\|^{2}}{{\mathbf{v}}^{\top}{\mathbf{B}}_{t}}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\chi^{2}(N-1).$
B.1. General results
Let ${{\mathbf{A}}}_{t}=(A_{t,1},\dots,A_{t,N})^{\top}$, with $t\geq 0$, be a multi-dimensional real stochastic processes, adapted to a filtration $(\mathcal{F}_{t})_{t}$, with the following dynamics:
(B.7) ${\mathbf{A}}_{t+1}={\mathbf{A}}_{t}-\frac{1}{t+1}(\phi^{*}I-\Phi^{\top}){\mathbf{A}}_{t}+\frac{1}{t+1}\Phi^{\top}\Delta{{\mathbf{M}}}_{t+1}+{\mathbf{R}}_{A,t+1}\$
where ${\mathbf{A}}_{0}$ is integrable and
  • $(i)$
    $\Phi^{\top}$ is a non-negative irreducible matrix with leading eigenvalue $0<\phi^{*}\leq 1$;
  • $(ii)$
[…]
Let ${\mathbf{u}}$ and ${\mathbf{v}}$ be the left and the right eigenvectors of $\Phi$ associated to $\phi^{*}$, with strictly positive components and such that ${\mathbf{v}}^{\top}{\boldsymbol{1}}=1\quad\mbox{and}\quad{\mathbf{v}}^{\top}{\mathbf{u}}=1.$ (Recall that $\phi^{*}$ is real and simple and it is possible to choose the components of these vectors all strictly positive because of the Frobenius-Perron theory). Set $\widetilde{A}_{t}={\mathbf{v}}^{\top}{\mathbf{A}}_{t}$.
Theorem B.1.
Under (i) and (ii) and assuming
  • (iii)
    $\sup_{t}E[\,|\widetilde{A}_{t}|\,]<+\infty$,
we have ${\widetilde{A}}_{t}\stackrel{{\scriptstyle\text{a.s.}}}{{\longrightarrow}}\widetilde{A}_{\infty}\,,$ where $\widetilde{A}_{\infty}$ is an integrable random variable.
Moreover, if we also assume
  • $(iv)$
    ${\mathbf{A}}_{0}$ square-integrable and $\sum_{t}w_{t}/(t+1)^{2}<+\infty$ where $w_{t}=\sum_{h=1}^{N}E[(\Delta M_{t+1,j})^{2}]$,
then we have $\sup_{t}E[(\widetilde{A}_{t})^{2}]<+\infty$ and so $\widetilde{A}_{t}$ converges to $\widetilde{A}_{\infty}$ also in quadratic mean (i.e. in $L^{2}$).
Note that (iii) is verified also when $(\widetilde{A}_{t})_{t}$ is uniformly integrable and in this case we also have that the convergence is in mean.
Proof.
Proof. By multiplying equation (B.7) by ${\mathbf{v}}^{\top}$ we obtain
(B.8) $\widetilde{A}_{t+1}=\widetilde{A}_{t}+\frac{1}{t+1}\phi^{*}\Delta\widetilde{M}_{t+1}+{\mathbf{v}}^{\top}{\mathbf{R}}_{A,t+1}\,,$
where $\Delta\widetilde{M}_{t+1}={\mathbf{v}}^{\top}\Delta{\mathbf{M}}_{t+1}$. Setting $\breve{M}_{t}=\sum_{n=1}^{t}\frac{1}{n}\Delta\widetilde{M}_{n}$ and $\widetilde{R}_{A,t+1}={\mathbf{v}}^{\top}{\mathbf{R}}_{A,t+1}=O(1/t^{1+\beta})$ (by (ii)), we have
(B.9) $\widetilde{A}_{t+1}-\widetilde{A}_{0}=\sum_{n=0}^{t}(\widetilde{A}_{n+1}-\widetilde{A}_{n})=\phi^{*}\breve{M}_{t+1}+\sum_{n=0}^{t}\widetilde{R}_{A,n+1}\,.$
Hence, since $\sup_{t}E[\,|\widetilde{A}_{t}|\,]<+\infty$ (by (iii)) and $\sum_{t}1/t^{1+\beta}<+\infty$, we also have $\sup_{t}E[\,|\breve{M}_{t}|\,]<+\infty$. Therefore, $(\breve{M}_{t})_{t}$ is a martingale bounded in $L^{1}$ and so it converges almost surely to an integrable random variable $\breve{M}_{\infty}$. It follows from (B.9) that $(\widetilde{A}_{t})$ converges almost surely to an integrable random variable $\widetilde{A}_{\infty}$. Moreover, we obtain
(B.10) $(\widetilde{A}_{t}-\widetilde{A}_{\infty})^{2}$ $=\big{(}\phi^{*}(\breve{M}_{t}-\breve{M}_{\infty})+\sum_{n\geq t}\widetilde{R}_{A,n+1}\big{)}^{2}$
$=(\phi^{*})^{2}(\breve{M}_{t}-\breve{M}_{\infty})^{2}+O(1/t^{\beta})\,.$
Now, we are going to prove that, under assumption (iv), we have $\sup_{t}(\widetilde{A}_{t})^{2}<+\infty$ so that (by (B.9)) $(\breve{M}_{t})_{t}$ is a martingale bounded in $L^{2}$ and so $\breve{M}_{\infty}$ is square-integrable and $\breve{M}_{t}$ converges in quadratic mean to it. By (B.10) this fact obviously implies that $\widetilde{A}_{t}$ converges in quadratic mean to $\widetilde{A}_{\infty}$.
We observe that, from (iv) and the fact that $(\Delta\widetilde{M}_{t+1})^{2}\leq C\sum_{j=1}^{N}(\Delta M_{t+1,j})^{2}$, we obtain
$E\Big{[}\,\Big{|}\frac{\Delta\widetilde{M}_{t+1}}{t+1}\widetilde{R}_{A,t+1}\Big{|}\,\Big{]}^{2}\leq E\Big{[}\,\frac{(\Delta\widetilde{M}_{t+1})^{2}}{(t+1)^{2}}\Big{]}E[(\widetilde{R}_{A,t+1})^{2}]\leq C\frac{w_{t}}{(t+1)^{2}}O(1/t^{2(1+\beta)})$
and so
$E\Big{[}\,\Big{|}\frac{\Delta\widetilde{M}_{t+1}}{t+1}\widetilde{R}_{A,t+1}\Big{|}\,\Big{]}=o(1/t^{1+\beta})\,.$
Therefore, from (B.8), since $\sup_{t}E[\widetilde{A}_{t}]<+\infty$, we get
$E[(\widetilde{A}_{t+1})^{2}]$ $\leq E[(\widetilde{A}_{t})^{2}]+(\phi^{*})^{2}C\frac{w_{t}}{(t+1)^{2}}+O\Big{(}\frac{1}{t^{2(1+\beta)}}\Big{)}+O\Big{(}\frac{\sup_{t}E[\widetilde{A}_{t}]}{t^{1+\beta}}\Big{)}+o\Big{(}\frac{1}{t^{1+\beta}}\Big{)}\,,$
$=E[(\widetilde{A}_{t})^{2}]+(\phi^{*})^{2}C\frac{w_{t}}{(t+1)^{2}}+O(1/t^{1+\beta})\,.$
Then, we find
$|E[(\widetilde{A}_{t})^{2}]-E[(\widetilde{A}_{0})^{2}]|$ $\leq\sum_{n=0}^{t-1}|E[(\widetilde{A}_{n+1})^{2}]-E[(\widetilde{A}_{n})^{2}]|$
$\leq(\phi^{*})^{2}C\sum_{n}\frac{w_{n}}{(n+1)^{2}}+\sum_{n}O(1/n^{1+\beta})<+\infty\,,$
where we have used (iv) in order to say that the first series is finite. Therefore, assuming ${\mathbf{A}}_{0}$ (and so $\widetilde{A}_{0}$) square-integrable, we have $\sup_{t}E[(\widetilde{A}_{t})^{2}]<+\infty$. ∎∎
Remark B.1 (non-negative case).
In particular, condition (iii) is verified when $\widetilde{A}_{t}$ is non-negative for each $t$. Indeed, if $\widetilde{A}_{t}$ is non-negative, we have, for each $t$,
$|E[\widetilde{A}_{t}]-E[\widetilde{A}_{0}]|\leq\sum_{n=0}^{t-1}|E[\widetilde{A}_{n+1}]-E[\widetilde{A}_{n}]|\leq\sum_{n}O(1/n^{1+\beta})$
and thus, since the last series is finite and $\widetilde{A}_{0}$ integrable, we have $\sup_{t}E[\widetilde{A}_{t}]<+\infty$, that is (iii) is verified.
Theorem B.2.
Assuming (i), (ii), (iii) and (iv), we have
${{\mathbf{A}}}_{t}\stackrel{{\scriptstyle\text{a.s.}/L^{2}}}{{\longrightarrow}}\widetilde{A}_{\infty}{{\mathbf{u}}}\,,$
where $\widetilde{A}_{\infty}$ is a square-integrable random variable.
Proof.
Proof. We firstly want to prove that we can neglect the term $R_{A,t+1}$ in the dynamics (B.7) of ${\mathbf{A}}_{t}$.
We recall that the matrix $\Phi^{\top}$ can be decomposed as
$\Phi^{\top}=\phi^{*}{\mathbf{u}}{\mathbf{v}}^{\top}+UDV^{\top}\,,$
where $D$ is the diagonal matrix whose elements are the eigenvalues of $\Phi$ (i.e. of $\Phi^{\top}$) different from $\phi^{*}$ and $U$ and $V$ denote the matrices whose columns are the left (right) and the right (left) eigenvectors of $\Phi$ (of $\Phi^{\top}$, respectively) associated to these eigenvalues, so that we have
(B.11) $V^{\top}{\mathbf{u}}=U^{\top}{\mathbf{v}}=0,\quad V^{\top}U=U^{\top}V=I,\quad I={\mathbf{u}}{\mathbf{v}}^{\top}+UV^{\top}\,.$
Therefore the dynamics of ${\mathbf{A}}_{t}$ can be rewritten as follows:
(B.12) ${\mathbf{A}}_{t+1}$ $=\Big{(}I-\frac{1}{t+1}U(I\phi^{*}-D)V^{\top}\Big{)}{\mathbf{A}}_{t}$
$\qquad+\frac{1}{t+1}\Phi^{\top}\Delta{{\mathbf{M}}}_{t+1}+{{\mathbf{R}}}_{A,t+1}.$
Let $\alpha_{j}=\phi^{*}-\phi_{j}$ with $\phi_{j}$ eigenvalue of $\Phi$ different from $\phi^{*}$. Then ${\mathcal{R}e}(\alpha_{j})>0$. Moreover, we have
(B.13) ${\mathbf{A}}_{t+1}$ $=\ C_{m_{0},t}{\mathbf{A}}_{m_{0}}+\sum_{k=m_{0}}^{t}\frac{1}{k+1}C_{k+1,t}\Phi^{\top}\Delta{\mathbf{M}}_{k+1}$
$\qquad+\sum_{k=m_{0}}^{t}C_{k+1,t}{\mathbf{R}}_{{\mathbf{A}},k+1}$
$=\ C_{m_{0},t}{\mathbf{A}}_{m_{0}}+\sum_{k=m_{0}}^{t}\frac{1}{k+1}C_{k+1,t}\Phi^{\top}\Delta{\mathbf{M}}_{k+1}+\boldsymbol{{\rho}}_{t+1}\,,$
where $m_{0}$ is such that ${\mathcal{R}e}(\alpha_{j})/(m_{0}+1)<1$ for each $j$ and
$C_{k+1,t}$ $=\prod_{m=k+1}^{t}\Big{(}I-\frac{1}{m+1}(I\phi^{*}-\Phi^{\top})\Big{)}=\prod_{m=k+1}^{t}\Big{(}I-\frac{1}{m+1}U(I\phi^{*}-D)V^{\top}\Big{)}$
$=U\bigg{(}\prod_{m=k+1}^{t}\Big{(}I-\frac{1}{m+1}(I\phi^{*}-D)\Big{)}\bigg{)}V^{\top},$
and so we can write $C_{k+1,t}=UA_{k+1,t}V^{\top}$ and
$A_{k+1,t}=\prod_{m=k+1}^{t}\Big{(}I-\frac{1}{m+1}(I\phi^{*}-D)\Big{)}.$
Moreover, setting for any $x\in{\mathbb{C}}$ with ${\mathcal{R}e}(x)/(m_{0}+1)<1$, $p_{m_{0}}(x)=1$ and $p_{k}(x)=\prod_{m=m_{0}}^{k}(1-\tfrac{x}{m+1})$ for $k\geq m_{0}$ and $F_{k+1,t}=\frac{p_{t}(x)}{p_{k}(x)}$ for $m_{0}-1\leq k\leq t-1$, from (Aletti et al. 2019, Lemma A.5) we get
$[A_{k+1,t}]_{jj}=F_{k+1,t}(\alpha_{j})\,.$
We now prove that $|\boldsymbol{{\rho}}_{t+1}|\stackrel{{\scriptstyle a.s.}}{{\to}}0$. To this end, first notice that $O(|C_{k+1,t}|)=O(|A_{k+1,t}|)$ and, setting $a_{2}^{*}={\mathcal{R}e}(\alpha_{2}^{*})=\phi^{*}-{\mathcal{R}e}(\phi^{*}_{2})$ with $\phi_{2}^{*}$ eigenvalue of $\Phi$, different form $\phi^{*}$, such that ${\mathcal{R}e}(\phi_{2}^{*})=\max_{j}\{{\mathcal{R}e}(\phi_{j})\}$, we have (see (Aletti et al. 2017, Lemma A.4))
$|A_{k+1,t}|=O\Big{(}\frac{|p_{t}(\alpha_{2}^{*})|}{|p_{k}(\alpha_{2}^{*})|}\Big{)}=O\Big{(}\Big{(}\frac{k}{t}\Big{)}^{a_{2}^{*}}\Big{)}\quad\mbox{for }k=m_{0},\dots,t-1,$
and simply $|A_{t+1,t}|=O(1)$ for $k=t$. Moreover, recalling that ${\mathbf{R}}_{A,t+1}=O(t^{-(1+\beta)}){\boldsymbol{1}}$ for some $\beta>0$, we have
$|\boldsymbol{{\rho}}_{t+1}|$ $=\Big{|}\sum_{k=m_{0}}^{t}C_{k+1,t}{\mathbf{R}}_{A,k+1}\Big{|}=O\Big{(}\sum_{k=m_{0}}^{t-1}\Big{(}\frac{k}{t}\Big{)}^{a_{2}^{*}}\frac{1}{k^{{1+\beta}}}\Big{)}+O(1/t^{1+\beta})$
$=O\Big{(}\frac{1}{t^{a_{2}^{*}}}\sum_{k=m_{0}}^{t-1}k^{a_{2}^{*}-1-\beta}\Big{)}+O(1/t^{1+\beta})\to 0,$
because $a_{2}^{*}>0$ and $\beta>0$.
Therefore, in all the sequel, without loss of generality, we can assume that ${\mathbf{A}}_{t}$ follows the dynamics (B.7) with ${\mathbf{R}}_{A,t+1}={\boldsymbol{0}}$.
We now decompose the vectorial process ${{\mathbf{A}}}_{t}$ by means of the Jordan representation of the matrix $\Phi$. Specifically, for any $\phi\in Sp(\Phi)\setminus\{\phi^{*}\}$, we can denote as $J_{\phi}$ the Jordan block and with $U_{\phi}$ and $V_{\phi}$ the matrices whose columns are, respectively, the left and right (possibly generalized) eigenvectors of $\Phi$ associated to the eigenvalue $\phi$, i.e.
$\Phi V_{\phi}=V_{\phi}J_{\phi}\qquad\text{and}\qquad U_{\phi}^{\top}\Phi=J_{\phi}U_{\phi}^{\top}.$
Then, we can consider the decomposition
${{\mathbf{A}}}_{t}=\widetilde{A}_{t}{{\mathbf{u}}}+\sum_{\phi\in Sp(\Phi)\setminus\{\phi^{*}\}}{{\mathbf{A}}}_{\phi,t}\,,$
where $\widetilde{A}_{t}={{\mathbf{v}}}^{\top}{{\mathbf{A}}}_{t}$ (as defined above) and ${{\mathbf{A}}}_{\phi,t}=U_{\phi}V_{\phi}^{\top}{{\mathbf{A}}}_{t}$.
We have already proven the almost sure convergence and the convergence in $L^{2}$ for $\widetilde{A}_{t}$ under (i), (ii), (iii) and (iv). In the following steps we are going to show that, under (i), (ii) and (iv), each ${{\mathbf{A}}}_{\phi,t}$ converges almost surely and in $L^{2}$ to zero. In particular, this last task will be done separately for the eigenvalues with $|\phi|<\phi^{*}$ and with $|\phi|=\phi^{*}$. Remember that the assumption that $\Phi$ (or, equivalently, $\Phi^{\top}$) is irreducible ensures that $\phi^{*}$ is real, simple and $|\phi|\leq\phi^{*}$ for any $\phi\in Sp(\Phi)$. Moreover, let us set $W_{t}=\sum_{h=1}^{N}E[(\Delta M_{t+1,j})^{2}\,|\,{\mathcal{F}}_{t}]$ and observe that assumption (iv) means $E[\sum_{t}W_{t}/(t+1)^{2}]=\sum_{t}E[W_{t}]/(t+1)^{2}=\sum_{t}w_{t}/(1+t)^{2}<+\infty$, which also implies $\sum_{t}W_{t}/(t+1)^{2}<+\infty$ almost surely.
Study of ${{\mathbf{A}}}_{\phi,t}$ with $|\phi|<\phi^{*}$.
Let $\breve{{\mathbf{A}}}_{t}=V_{\phi}^{\top}{\mathbf{A}}_{t}$ and since ${{\mathbf{A}}}_{\phi,t}=U_{\phi}V_{\phi}^{\top}{\mathbf{A}}_{t}=U_{\phi}\breve{{\mathbf{A}}}_{t}$, it is enough to prove that $\|\breve{{\mathbf{A}}}_{t}\|^{2}$ converges a.s. and in $L^{2}$ to zero. To this end, by multiplying equation (B.7) by $V_{\phi}^{\top}$, we have
$\breve{{\mathbf{A}}}_{t+1}=\Big{(}I-\frac{1}{t+1}(\phi^{*}I-J_{\phi}^{\top})\Big{)}\breve{{\mathbf{A}}}_{t}+\frac{1}{t+1}J_{\phi}^{\top}V_{\phi}^{\top}\Delta{{\mathbf{M}}}_{t+1}.$
Then, since for any real matrix $Q$ we can write
(B.14) $E[\Delta{\mathbf{M}}_{t+1}^{\top}Q\Delta{\mathbf{M}}_{t+1}|\mathcal{F}_{t}]$ $=\sum_{j=1}^{N}q_{jj}^{2}E[(\Delta M_{j,t+1})^{2}|\mathcal{F}_{t}]$
$\leq\max_{j}{q_{jj}^{2}}W_{t},$
we have that
$E[\|\breve{{\mathbf{A}}}_{t+1}\|^{2}$ $|\mathcal{F}_{t}]$
$\leq\big{\|}\bigg{(}\Big{(}1-\frac{\phi^{*}}{t+1}\Big{)}I+\frac{1}{t+1}J_{\phi}\bigg{)}\breve{{\mathbf{A}}}_{t}\big{\|}^{2}+\frac{1}{(t+1)^{2}}\sum_{j=1}^{N}[\bar{V}_{\phi}\bar{J}_{\phi}J_{\phi}^{\top}V_{\phi}^{\top}]_{jj}^{2}E[(\Delta M_{j,t+1})^{2}|\mathcal{F}_{t}]$
$\leq\Big{(}1-\frac{\phi^{*}}{t+1}+\frac{\|J_{\phi}\|_{2,2}}{t+1}\Big{)}^{2}\|\breve{{\mathbf{A}}}_{t}\|^{2}+\frac{1}{(t+1)^{2}}\max_{j}\{[\bar{V}_{\phi}\bar{J}_{\phi}J_{\phi}^{\top}V_{\phi}^{\top}]^{2}_{jj}\}W_{t}.$
Then, regarding the first term, we note that
$\Big{(}1-\frac{\phi^{*}}{t+1}+\frac{\|J_{\phi}\|_{2,2}}{t+1}\Big{)}^{2}\leq\Big{(}1-\frac{\phi^{*}}{t+1}+\frac{|\phi|+\phi^{*}}{2(t+1)}\Big{)}^{2}=\Big{(}1-\frac{\phi^{*}-|\phi|}{2(t+1)}\Big{)}^{2},$
and so
$E[\|\breve{{\mathbf{A}}}_{t+1}\|^{2}|\mathcal{F}_{t}]\leq\Big{(}1-\frac{\phi^{*}-|\phi|}{2(t+1)}\Big{)}^{2}\|\breve{{\mathbf{A}}}_{t}\|^{2}+\frac{C}{(t+1)^{2}}W_{t}.$
Therefore, since $\phi^{*}>|\phi|$ and by (iv), the process $\|\breve{{\mathbf{A}}}_{t}\|^{2}$ is a non-negative almost supermartingale so that it converges almost surely (see Robbins and Siegmund (1971)). Moreover, by applying the expectation, we obtain
$E[\|\breve{{\mathbf{A}}}_{t+1}\|^{2}]\leq\Big{(}1-\frac{\phi^{*}-|\phi|}{2(t+1)}\Big{)}^{2}E[\|\breve{{\mathbf{A}}}_{t}\|^{2}]+\frac{C}{(t+1)^{2}}E[W_{t}],$
which, since $\sum_{t}(\phi^{*}-|\phi|)/(t+1)=+\infty$, by (iv) and (Aletti et al. 2023a, Lemma S1.6), we can conclude that $\|\breve{{\mathbf{A}}}_{t}\|\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}0$, and hence $\breve{{\mathbf{A}}}_{t}\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}{\boldsymbol{0}}$.
Study of ${{\mathbf{A}}}_{\phi,t}$ with $|\phi|=\phi^{*}$.
From the Frobenius-Perron theory, we know that each eigenvalue with maximum modulus is simple. Let us denote the corresponding right and left eigenvectors by ${\mathbf{v}}_{\phi}$ and ${\mathbf{u}}_{\phi}$ of $\Phi$. Then, set $a_{\phi,t}={\mathbf{v}}_{\phi}^{\top}{\mathbf{A}}_{t}$ so that, since we have ${{\mathbf{A}}}_{\phi,t}={\mathbf{u}}_{\phi}{\mathbf{v}}_{\phi}^{\top}{\mathbf{A}}_{t}={\mathbf{u}}_{\phi}a_{\phi,t}$, it is enough to prove that $|a_{\phi,t}|$ converges to zero almost surely and in $L^{2}$. To this end, by multiplying equation (B.7) by ${\mathbf{v}}_{\phi}^{\top}$, we have
${a}_{\phi,t+1}=\Big{(}1-\frac{1}{t+1}(\phi^{*}-\phi)\Big{)}{a}_{\phi,t}+\frac{\phi}{t+1}{\mathbf{v}}_{\phi}^{\top}\Delta{{\mathbf{M}}}_{t+1}.$
Then, using (B.14), we have that
$E[|a_{\phi,t+1}|^{2}|\mathcal{F}_{t}]$ $\leq\Big{|}1-\frac{\phi^{*}}{t+1}+\frac{\phi}{t+1}\Big{|}^{2}|a_{\phi,t}|^{2}+\frac{|\phi|^{2}}{(t+1)^{2}}\sum_{j=1}^{N}|v_{j}|^{2}E[(\Delta M_{j,t+1})^{2}|\mathcal{F}_{t}]$
$\leq\Big{|}1-\frac{\phi^{*}}{t+1}+\frac{\phi}{t+1}\Big{|}^{2}|a_{\phi,t}|^{2}+\frac{|\phi|^{2}}{(t+1)^{2}}\max_{j}\{|v_{j}|^{2}\}W_{t}.$
Then, regarding the first term we have that
$\Big{|}1-\frac{\phi^{*}}{t+1}+\frac{\phi}{t+1}\Big{|}^{2}$ $=\Big{(}1-\frac{\phi^{*}}{t+1}+\frac{\mathcal{R}e(\phi)}{t+1}\Big{)}^{2}+\Big{(}\frac{\mathcal{I}m(\phi)}{t+1}\Big{)}^{2}$
$=1+\Big{(}\frac{\phi^{*}-\mathcal{R}e(\phi)}{t+1}\Big{)}^{2}-2\Big{(}\frac{\phi^{*}-\mathcal{R}e(\phi)}{t+1}\Big{)}+\Big{(}\frac{\mathcal{I}m(\phi)}{t+1}\Big{)}^{2}$
$=1-\Big{(}\frac{2(\phi^{*}-\mathcal{R}e(\phi))}{t+1}\Big{)}+\Big{(}\frac{\phi^{*2}-2\phi^{*}\mathcal{R}e(\phi)+\mathcal{R}e(\phi)^{2}+\mathcal{I}m(\phi)^{2}}{(t+1)^{2}}\Big{)}$
$=1-\Big{(}\frac{2(\phi^{*}-\mathcal{R}e(\phi))}{t+1}\Big{)}+\Big{(}\frac{2\phi^{*}(\phi^{*}-\mathcal{R}e(\phi))}{(t+1)^{2}}\Big{)}$
$=1-2\Big{(}\frac{1}{t+1}-\frac{\phi^{*}}{(t+1)^{2}}\Big{)}(\phi^{*}-\mathcal{R}e(\phi))$
and so
$E[|a_{\phi,t+1}|^{2}|\mathcal{F}_{t}]$ $\leq\Big{(}1-2\Big{(}\frac{1}{t+1}-\frac{\phi^{*}}{(t+1)^{2}}\Big{)}(\phi^{*}-\mathcal{R}e(\phi))\Big{)}|a_{\phi,t}|^{2}+\frac{C}{(t+1)^{2}}W_{t}.$
Therefore, since $\phi^{*}>\mathcal{R}e(\phi)$ and by (iv), the process $|a_{\phi,t}|^{2}$ is a non-negative almost supermartingale so that it converges almost surely (see Robbins and Siegmund (1971)). Moreover, by applying the expectation, we obtain
$E\big{[}|a_{\phi,t+1}|^{2}\big{]}$ $\leq\Big{(}1-2\Big{(}\frac{1}{t+1}-\frac{\phi^{*}}{(t+1)^{2}}\Big{)}(\phi^{*}-\mathcal{R}e(\phi))\Big{)}E\big{[}|a_{\phi,t}|^{2}\big{]}+\frac{C}{(t+1)^{2}}E[W_{t}]\,.$
Since $\sum_{t}(1/(t+1)-\phi^{*}/(t+1)^{2})=+\infty$ and by (iv) and (Aletti et al. 2023a, Lemma S1.6), we can conclude that $|a_{\phi,t}|\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}0$, and hence $a_{\phi,t}\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}0$. ∎∎
B.2. Technical result
Let ${\mathcal{W}}=(\mathcal{W}_{t})_{t\geq 0}$ be a bounded stochastic process with the following dynamics
$\mathcal{W}_{t+1}=\Big{(}1-\frac{1}{t+1}\Big{)}\mathcal{W}_{t}+\frac{1}{t+1}Y_{t+1}+R_{t+1},\quad t\geq 0,$
where $(Y_{t})_{t}$ is a bounded stochastic process and $R_{t}=O(1/t^{1+\beta})$ with $\beta>0$. Define a bounded stochastic process $\mathcal{V}=(\mathcal{V}_{t})_{t\geq 0}$ with dynamics
$\mathcal{V}_{t+1}$ $=\Big{(}1-\frac{1}{t+2}\Big{)}\mathcal{V}_{t}+\frac{1}{t+2}Y_{t+1},\quad t\geq 0\,.$
Then
$|\mathcal{W}_{t}-\mathcal{V}_{t}|=O(1/t)|\mathcal{W}_{0}-\mathcal{V}_{0}|+\frac{1}{t}\sum_{n=0}^{t}O(n^{-\beta^{\prime}})\,,$
where $\beta^{\prime}=\min(\beta,1)$.
Indeed, we can observe that we can write
$\mathcal{W}_{t+1}=\Big{(}1-\frac{1}{t+2}\Big{)}\mathcal{W}_{t}+\frac{1}{t+2}Y_{t+1}+R^{\prime}_{t+1},\quad t\geq 0,$
with $R^{\prime}_{t+1}=R_{t+1}+O(1/t^{2})=O(1/t^{1+\beta^{\prime}})$ and so
$\mathcal{W}_{t+1}=C_{0,t}\mathcal{W}_{0}+\sum_{n=0}^{t}C_{n+1,t}\frac{Y_{n+1}}{n+2}+\sum_{n=0}^{t}C_{n+1,t}R^{\prime}_{n+1}$
where
$C_{0,t}$ $=\prod_{m=0}^{t}\Big{(}1-\frac{1}{m+2}\Big{)}=O(1/t)$
$C_{n+1,t}$ $=\prod_{m=n+1}^{t}\Big{(}1-\frac{1}{m+2}\Big{)}=\frac{\prod_{m=0}^{t}\big{(}1-\frac{1}{m+2}\big{)}}{\prod_{m=0}^{n}\big{(}1-\frac{1}{m+2}\big{)}}=O(n/t)$
Similarly, we have
$\mathcal{V}_{t+1}=C_{0,t}\mathcal{V}_{0}+\sum_{n=0}^{t}C_{n+1,t}\frac{Y_{n+1}}{n+2}.$
Hence, we obtain
$|\mathcal{W}_{t+1}-\mathcal{V}_{t+1}|$ $\leq C_{0,t}|\mathcal{W}_{0}-\mathcal{V}_{0}|+\Big{|}\sum_{n=0}^{t}C_{n+1,t}R^{\prime}_{n+1}\Big{|}$
$=O(1/t)|\mathcal{W}_{0}-\mathcal{V}_{0}|+\sum_{n=0}^{t}O\Big{(}\frac{n}{t}\frac{1}{n^{1+\beta^{\prime}}}\Big{)}$
$=O(1/t)|\mathcal{W}_{0}-\mathcal{V}_{0}|+\frac{1}{t}\sum_{n=0}^{t}O(n^{-\beta^{\prime}})\,.$
Appendix C The case of a reducible interaction matrix
When the interaction matrix $\Gamma$ is not irreducible, it is possible to decompose it in irreducible sub-matrices such that the union of the spectra of the sub-matrices coincides with the spectrum of the original matrix. In the following, we will describe an heuristic argument (also employed in Iacopini et al. (2020) and Aletti et al. (2023a)), useful in order to detect the rate at which each $S_{t,h}$ grows along time in the case of a general matrix $\Gamma$.
The dynamics that rules the vectorial process ${\mathbf{S}}_{t}=(S_{t,1},\dots,S_{t,N})^{\top}$ can be approximated (as $t\to+\infty$) by the linear system of (deterministic) differential equations
$\dot{\mathbf{s}}(t)=\Gamma\frac{\mathbf{s}(t)}{t}$
and hence we can say that ${\mathbf{S}}_{t}\approx{\mathbf{s}}(t)$ for $t\to+\infty$. By the change of variable $t=e^{z}$, we get
$\dot{{\mathbf{s}}}(z)=\Gamma{\mathbf{s}}(z)\,,$
whose general solution is given by ${\mathbf{s}}(z)=e^{\Gamma z}\mathbf{c}$. Now, the term $e^{\Gamma z}$ can be expressed using the canonical Jordan form of the matrix $\Gamma$, so that we obtain
${\mathbf{s}}(z)=\sum_{k=1}^{r}e^{\gamma_{k}z}\sum_{i=0}^{p_{k}-1}z^{i}\mathbf{c}_{i},$
where $\gamma_{1},\dots,\gamma_{r}$ are the distinct eigenvalues of $\Gamma$, $p_{1},\dots,p_{r}$ are the sizes of the corresponding Jordan blocks and $\mathbf{c}_{i}$ are suitable vectors related to $\mathbf{c}$ and to the generalized eigenvectors of $\Gamma$. Indeed, we can write $\Gamma$ as $PJP^{-1}$, where $J$ is its canonical Jordan form and $P$ is a suitable invertible matrix of generalized eigenvectors. Therefore, we have $e^{\Gamma z}=Pe^{Jz}P^{-1}$, where $e^{Jz}$ is a block matrix with blocks of the form $e^{J_{k}z}$ with $J_{k}$ block in $J$. On the other hand, if $J_{k}=\gamma_{k}I+N_{k}$ is a generic Jordan block of $\Gamma$ with size $p_{k}$ and associated to the eigenvalue $\gamma_{k}$, we have
$e^{J_{k}z}=e^{\gamma_{k}z}e^{N_{k}z}=e^{\gamma_{k}z}\sum_{i=0}^{p_{k}-1}\frac{z^{i}}{(i-1)!}N_{k}^{i}\,.$
Changing the variable from $z$ to $t$, we find
(C.1) ${\mathbf{S}}_{t}\approx{\mathbf{s}}(t)=\sum_{k=1}^{r}t^{\gamma_{k}}\sum_{i=0}^{p_{k}-1}\ln^{i}(t)\mathbf{c}_{i}$
and so the rate at which $S_{t,h}$ increases is given by the leading term in the expression of $s_{h}(t)$.
Appendix D Choice of the threshold (robustness check)
Figure 6. Behavior of the mean index per category over the years.
In this section we discuss the choice of the threshold $\tau$ that has been used in the data analysis of Section 4 to define when a given patent $n$ can be considered a success for (or in) category $h$, that is, if and only if $I_{n,h}>\tau$. Since we want to identify as a “success” only patents that have an extraordinary impact on at least one category, the value of the threshold $\tau$ should be greater than the index value of the vast majority of patents in the data set. However, this requirement is not so stringent in the sense that the percentage of patents with $(I_{n,h}>\tau)$ is already around $1\%$ for $\tau=0.1$ and is below $0.1\%$ for any $\tau>0.5$ (see Figure 6 and Table 3 for further details). This means that the majority of patents in the data set have an index value very low (precisely, below $0.1$). This is in accordance with Squicciarini et al. (2013), who observed that only a very small subset of patents typically receives a large number of forward citations and the mean value of the forward index decreases along time.
category $(I_{n,h}>0.1)\%$ $(I_{n,h}>0.3)\%$ $(I_{n,h}>0.5)\%$ $(I_{n,h}>0.7)\%$ $(I_{n,h}>0.9)\%$
A 1.640 % 0.185 % 0.045 % 0.017 % 0.008 %
B 2.440 % 0.250 % 0.059 % 0.021 % 0.008 %
C 1.180 % 0.154 % 0.042 % 0.017 % 0.008 %
D 0.554 % 0.1410 % 0.031 % 0.017 % 0.011 %
E 1.070 % 0.149 % 0.038 % 0.017 % 0.009 %
F 1.420 % 0.170 % 0.043 % 0.018 % 0.008 %
G 1.450 % 0.153 % 0.038 % 0.016 % 0.008 %
H 1.350 % 0.150 % 0.037 % 0.016 % 0.008 %
Table 3. For each category $h$, percentage of patents with a value of the index $I_{n,h}$ greater than $\tau=0.1,0.3,0.5,0.7,0.9$.
Figure 7. Different slopes estimated through linear regression on $S_{t,h}$, in $\log_{10}-\log_{10}$ scale, for each category $h$ (color line) and each value of the threshold $\tau$ (x-axis). The solid black line indicates the common slope $\widehat{\gamma}^{*}$ estimated imposing the same slope for all categories. The dashed horizontal black line indicates the common slope $\widehat{\gamma}^{*}=0.689$ estimated for $\tau=0.8$ in Section 4.
Then, to verify the robustness of the results of the paper with respect to the choice of the threshold $\tau$, we performed the data analysis presented in Section 4 multiple times, each one using a different threshold value. In particular, we have performed a linear regression for every process $S_{t,h}$, in the $\log_{10}-\log_{10}$ scale, allowing the slopes to be different between categories and also imposing a common slope. The results are collected in Figure 7, where we can see that the variability presented by the slopes is essentially very similar for any value of $\tau>0.2$. In addition, when we perform the linear regression that imposes the same slope for all categories, the estimated common slope $\widehat{\gamma}^{*}$ is always very close to the value $0.689$, that is the one estimated with $\tau=0.8$ in Section 4. Finally, we have calculated the goodness-of-fit index $R^{2}$ obtained imposing a common slope and the one obtained allowing the slopes to be different across categories, and they are always very close to each other and always higher than 0.95 for any value of $\tau$ (see Figure 8).
Figure 8. Goodness-of-fit index $R^{2}$ of the linear regression on $S_{t,h}$, in $\log_{10}-\log_{10}$ scale, for each value of the threshold $\tau$ (x-axis). The red line indicates the $R^{2}$ obtained imposing a common slope and the blue line the $R^{2}$ obtained allowing the slopes to be different across categories. Both the lines are always very close to each other and always higher than $0.95$ for any value of $\tau$.